### v1

# Self-NPO: Negative Preference Optimization of Diffusion Models by Simply Learning from Itself without Explicit Preference Annotations

Abstract
Diffusion models have demonstrated remarkable success in various visual generation tasks, including image, video, and 3D content generation. Preference optimization (PO) is a prominent and growing area of research that aims to align these models with human preferences. While existing PO methods primarily concentrate on producing favorable outputs, they often overlook the significance of classifier-free guidance (CFG) in mitigating undesirable results. Diffusion-NPO addresses this gap by introducing negative preference optimization (NPO), training models to generate outputs opposite to human preferences and thereby steering them away from unfavorable outcomes. However, prior NPO approaches, including Diffusion-NPO, rely on costly and fragile procedures for obtaining explicit preference annotations (e.g., manual pairwise labeling or reward model training), limiting their practicality in domains where such data are scarce or difficult to acquire. In this work, we introduce Self-NPO, a Negative Preference Optimization approach that learns exclusively from the model itself, thereby eliminating the need for manual data labeling or reward model training. Moreover, our method is highly efficient and does not require exhaustive data sampling. We demonstrate that Self-NPO integrates seamlessly into widely used diffusion models, including SD1.5, SDXL, and CogVideoX, as well as models already optimized for human preferences, consistently enhancing both their generation quality and alignment with human preferences.
1 Introduction
Over the past few years, diffusion models ho2020denoising ; song2020denoising ; song2020score have made significant strides in various visual generation tasks, including image karras2022elucidating ; rombach2022high ; wang2024rectified ; podell2023sdxl ; diffusionbeatgan ; meng2021sdedit ; teng2024dim ; ma2024storynizor ; li2024ecnet ; sun2024journeydb , video blattmann2023align ; mao2024osv ; shi2024motion ; makeavideo ; chen2024videocrafter2 ; bian2025gs ; NEURIPS2022_f7f47a73 ; wang2024animatelcm ; wang2025your, and 3D generation gao2024cat3d ; poole2022dreamfusion ; li2025connecting ; yan2025consistent ; sun2024genca ; lai2025unleashing. However, diffusion models trained on massive unfiltered data (e.g., image-text pairs) often generate results that do not align well with human preferences. The growing interest in aligning these models for human preference-aligned generation has led to the development of preference optimization (PO) methods. In general, current PO methods can be categorized into four types:
Figure 1: Visual Comparison. Prompts from left to right, from top to bottom are “galaxy cat”, “cat as a mafia boss”, and “A real picture of an Canguru with sunglasses at the sunset holding a stop sign”.
Differentiable Reward (DR): In this approach, images generated via iterative denoising are evaluated by a pre-trained reward model, and the diffusion models are subsequently refined through gradient-based backpropagation to align with the reward model xu2024imagereward ; prabhudesai2024video ; zhang2024large ; wu2023hpsv1 ; wu2023human ; wu2024deepreward .
Reinforcement Learning (RL): Here, the denoising process in diffusion models is treated as a Markov decision process (MDP) puterman2014markov ; sutton2018reinforcement . Techniques like PPO and its variants are employed to optimize preferences schulman2017proximal . Images are generated and evaluated dynamically based on reward feedback, with the goal of maximizing the chances of producing images that yield higher rewards.
Direct Preference Optimization (DPO): DPO streamlines training by reframing the reinforcement learning objective into a simulation-free framework rafailov2024direct ; wallace2024diffusion , eliminating the need for separate reward models or online image generation and evaluation. Instead, it relies on fine-tuning with existing datasets of paired preference annotations. Some works yuan2024self ; deng2024enhancing ; gu2024diffusion ; zhang2024seppo construct preference pairs by using high-quality demonstration data as positive examples and self-generated data as negative examples. We have more discussion with these works as they also mention the idea of self-generated data.
Negative Preference Optimization (Diffusion-NPO): Despite previous efforts to make models generate human-aligned images, Diffusion-NPO wang2025diffusionnpo firstly raised an important question: How can a model know to avoid generating poor images if it only knows how to generate good ones without understanding what is bad? Diffusion-NPO introduced the concept of negative preference optimization (NPO), which trains models to generate outputs that oppose human preferences, effectively improving performance by guiding the model away from undesired results though the classifier-free guidance (CFG) ho2022classifier ; karras2024guiding ; Shen_2024_CVPR ; ahn2024self .
However, we argue that a major limitation of existing methods including DR, RL, DPO, and Diffusion-NPO, is their reliance on explicit preference annotations, such as the expensive manual collection of preference data pairs kirstain2023pick ; wu2023hpsv2 and the challenging and fragile process of reward model training xu2024imagereward ; wu2023hpsv2 ; schuhmann2022laion . This heavy dependency on preference annotations makes these approaches less practical, particularly in domains where obtaining or labeling preference data is expensive or infeasible.
In this work, we introduce Self-NPO, a novel approach for negative preference optimization (NPO) that can be done in an explicit preference annotation-free manner. Specifically, our training objective is to apply negative preference optimization to a pre-trained model, making its predictions more aligned with the opposite direction of human preferences. This facilitates unconditional/negative-conditional outputs in CFG, thus reducing the probability of generating results that do not align with human preferences.
To implement Self-NPO, we have three key observations:
  1. 1.
    Controlled weakening of generative ability. People typically do not prefer corrupted generated images. For example, blurry details or structural errors make an image less appealing. Therefore, weakening the generative ability of a pre-trained diffusion model is, to some extent, equivalent to negative preference optimization. However, this weakening of generative ability cannot be arbitrary. In classifier-free guidance, we typically have the equation
    $\boldsymbol{\epsilon}^{\omega}=(\omega+1)\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{pos}}(\mathbf{x}_{t},t,\boldsymbol{c})-\omega\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{neg}}(\mathbf{x}_{t},t,\boldsymbol{c^{\prime}})$ (1)
    To avoid altering the variance of the epsilon predictions, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{neg}}(\mathbf{x}_{t},t,\boldsymbol{c^{\prime}})$ should maintain a certain level of correlation with $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{pos}}(\mathbf{x}_{t},t,\boldsymbol{c})$. For example, for two completely independent Gaussian noises, the above operation would result in a variance of $12\omega^{2}+2\omega+$ 1.
  2. 2.
    Weakening via self-generated data. This weakening, which satisfies the above conditions, can be achieved by learning from the model’s own generated data, which can be regarded as a target distribution regularized weakening.
  3. 3.
    Efficient tuning without full diffusion. We do not need to fully execute iterative diffusion generation process to implement self-learning. We propose a novel tuning strategy, which allows us to fine tune diffusion models on the partially excuted iterative diffusion generation results. This avoids the huge generation costs of preparing data for self-learning.
We demonstrate that Self-NPO can be seamlessly integrated with existing models, such as SD1.5 rombach2022high , SDXL podell2023sdxl , and CogVideoX yang2024cogvideox , as well as models previously fine-tuned through preference optimization, consistently improving their alignment with human preferences.
Figure 2: Motivation of Self-NPO. Preference Optimization (PO) optimizes the preference aligned model by training on preference annotations. Negative Preference Optimization (NPO) optimizes the negative preference aligned model by training on switched preference annotations. Self-NPO optimizes the negative preference aligned model by truncated diffusion fine-tuning on truncated self-generated data.
2 Preliminary
Preliminary of CFG. CFG has become a necessary and important technique for improving generation quality and text alignment of diffusion models. For convenience, we focus our discussion on the general formal of diffusion models, i.e., $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ kingma2021variational . Suppose we learn a score estimator from a epsilon prediction neural network $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c},t)$, and we have $\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)=-\frac{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)}{\sigma_{t}}$. The sample prediction at timestep $t$ of the score estimator is formulated as
$\hat{\boldsymbol{x}}_{0}=\frac{1}{\alpha_{t}}(\boldsymbol{x}_{t}+\sigma^{2}\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t))\,.$ (2)
Applying the CFG is equivalent to add an additional score term, that is, we replace $\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)$ in Eq. 2 with the following term,
$\begin{split}\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)+\nabla_{\boldsymbol{x}_{t}}\log\left[\frac{\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)}{\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c}^{\prime};t)}\right]^{\omega}\,,\end{split}$ (3)
where $\omega$ is to control the strength of CFG, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are conditional and unconditional/negative-conditional inputs, respectively. It is apparent that the generation will be pushed to high probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)$ and relatively low probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c}^{\prime};t)$. Write the above equation into the epsilon format, and then we have
$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\theta}}=(\omega+1){\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c},t)}-\omega{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c}^{\prime},t)}\,.$ (4)
Diffusion-NPO wang2025diffusionnpo . Diffusion models use classifier-free guidance (CFG) to improve generation quality by combining conditional and negative-conditional outputs to favor preferred results. Traditional preference optimization aligns outputs with human preferences but overlooks avoiding undesirable outputs. Negative Preference Optimization (NPO), introduced in Diffusion-NPO, trains an additional model to counteract human preferences, reducing undesired outputs via CFG.
Training with Diffusion-NPO. NPO adapts existing preference optimization methods without new datasets or strategies. For reward-based methods, the negative preference reward is:
$R_{\text{NPO}}(\mathbf{x},\boldsymbol{c})=1-R(\mathbf{x},\boldsymbol{c}),$ (5)
where $R(\mathbf{x},\boldsymbol{c})\in[0,1]$ is a reward model (e.g., HPSv2). For DPO-based methods, preference pairs $r=(\mathbf{x}_{0},\mathbf{x}_{1},\boldsymbol{c})$, with $\mathbf{x}_{1}$ preferred, are reversed:
$r_{\text{NPO}}=(\mathbf{x}_{1},\mathbf{x}_{0},\boldsymbol{c}).$ (6)
Inference with Diffusion-NPO. Let $\boldsymbol{\theta}$ denote the base model weights, with $\boldsymbol{\eta}$ and $\boldsymbol{\delta}$ representing the weight offsets after positive and negative preference optimization, respectively. The preference optimized model and negative preference optimized models can be represented as: $\boldsymbol{\theta}_{pos}=\boldsymbol{\theta}+\boldsymbol{\eta},\quad\boldsymbol{\theta}_{neg}=\boldsymbol{\theta}+\boldsymbol{\delta}.$ Classifier-free guidance is then applied as: $\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\theta}}=(\omega+1){\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{pos}}}-\omega{\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{neg}}}.$ However, this often results in significant output discrepancies. To mitigate this, Diffusion-NPO modifies the negative weights to include a combination of the positive and negative offsets:
$\boldsymbol{\theta}_{neg}=\boldsymbol{\theta}+\alpha\boldsymbol{\eta}+\beta\boldsymbol{\delta},\quad\alpha,\beta\in[0,1],$ (7)
which ensures more stable and correlated outputs.
3 Methodology
Figure 3: Visual comparison on Stable Diffusion XL. Prompts: “A cute blue cat.”, “An attractive young woman rolling her eyes.”, and “Close-up of a head with smoke from ears, watching a smartphone, dynamic angle”.
3.1 NPO via Self-Generated Data
Reinforcement Learning with Human Feedback (RLHF) rafailov2024direct ; wallace2024diffusion ; griffith2013policy ; liu2025improving aims to optimize a conditional distribution $\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})$, where $\boldsymbol{c}\sim\mathcal{D}_{c}$, in order to maximize the expectation value of the associated reward model $R(\boldsymbol{c},\mathbf{x}_{0})$. Simultaneously, the optimization imposes a regularization term that penalizes deviations from a reference distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})$. Formally, the objective function is formulated as follows:
$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R(\boldsymbol{c},\mathbf{x}_{0})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,,\end{split}$ (8)
where $\beta$ is a hyper-parameter that governs the balance between the reward maximization and the regularization effect imposed by the KL-divergence. The regularization is crucial because it prevents the model from straying too far from the distribution where the reward model is reliable, while also preserving generation diversity and preventing mode collapse to a single high-reward sample. Replacing $R(\mathbf{x}_{0},\boldsymbol{c})$ with $R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})$ yields the RLHF objective for Negative Preference Optimization (NPO):
$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R_{\text{NPO}}(\boldsymbol{c},\mathbf{x}_{0})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,.\end{split}$ (9)
where the first term encourages the samples generated from $\mathbb{P}_{\boldsymbol{\theta}}$ to have low reward scores, while the second term constrains the learned distribution to remain close to $\mathbb{P}_{\text{ref}}$.
Observing that a diffusion model can naturally produce undesirable outputs (e.g., disordered compositions, incoherent structures, blurry details), such low-quality samples inherently yield low reward scores. Consequently, exploiting self-generated data provides a straightforward path toward distribution-regularized negative preference optimization. Figure 2 summarizes the core idea and situates it among related methods. This approach is justified on two grounds: 1) Distribution preservation. Learning from self-generated data preserves the original distributional properties of the model. 2) Reward reduction. Self-generated samples exhibit low reward scores for several reasons: a) Hallucination. Generation of unrealistic objects, artifacts, or contextually inconsistent outputs, which result in low reward scores due to poor quality or implausibility of the generated content. b) Mode collapse. The model may fail to generate a diverse set of outputs, leading to a lack of coverage across the entire data distribution and, consequently, lower reward scores due to insufficient variety. c) Optimization. Imperfect optimization within the diffusion model itself can produce suboptimal results, such as blurry or noisy details, which leads to lower reward scores due to the misalignment between generated outputs and the desired target distribution. Indeed, we find it is helpful to intentionally corrupt the ODE sampling process to some extend, thus increasing the probability of reward reduction. We have more discussion in the Supplementary IV.1.
Figure 4: Truncated Diffusion Fine-Tuning. Baseline method requires full simulation of generation process, introducing huge amount of generation costs. Naive Truncation reduces the simulation costs but introduces distribution mismatch. Truncated Diffusion Fine-tuning reduces the simulation costs and maintains the target distribution.
However, despite its conceptual simplicity, fine-tuning a diffusion model on self-generated data can be computationally expensive, as it requires generating large amounts of training samples. To address this, we introduce a truncated diffusion fine-tuning strategy. This method updates the model using partially generated diffusion samples, thereby avoiding the high costs of performing full diffusion simulations during training.
3.2 Truncated Diffusion Fine-tuning
Baseline: Fine-tuning on fully generated data. We denote the reference diffusion model for data generation as $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})$, which models the adjacent timesteps transition conditional distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t-1}^{ref}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})$ for $t=1,2,\dots,T$. We can achieve data sampling by iteratively call the transition distribution,
$\mathbf{x}^{ref}_{T}\rightarrow\mathbf{x}^{ref}_{T-1}\rightarrow\dots\mathbf{x}^{ref}_{t}\rightarrow\dots\mathbf{x}^{ref}_{1}\rightarrow\mathbf{x}^{ref}_{0}\,,$ (10)
where $\mathbf{x}^{ref}_{T}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$. After obtaining the $\mathbf{x}_{t}^{ref}$, the learning of $\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{0}\mid\boldsymbol{c})$ is equivalent to the learning a $\mathbf{x}_{0}$-prediction neural network $\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t,\boldsymbol{c})$ with standard diffusion loss following previous work:
$\min\limits_{\boldsymbol{\theta}}\mathbb{E}_{t,\boldsymbol{\epsilon}}\|\mathbf{x}_{0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}((\mathbf{x}^{ref}_{0})_{t},t,\boldsymbol{c})\|_{2}^{2},$ (11)
where $t$ is sampled from $\{1,2,\dots,T\}$, and $(\mathbf{x}^{ref}_{0})_{t}=\alpha_{t}\mathbf{x}^{ref}_{0}+\sigma_{t}\boldsymbol{\epsilon}$.
Our method: Truncated Diffusion Fine-tuning. According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$, based on the conditional $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}_{t}^{ref})=\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}|\mathbf{x}_{0})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}$ according to Bayes’ rule, is related to the score of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ efron2011tweedie ; zhou2024score . Specifically, the expectation is:
$\mathbb{E}_{\mathbf{x}_{0}\sim\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}^{ref}_{t})}[\mathbf{x}_{0}|\mathbf{x}_{t}^{ref}]=\mathbf{x}_{t}+\sigma^{2}_{t}\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c}).$ (12)
In this equation, the expectation $\mathbb{E}_{\mathbf{x}_{0}\sim\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}^{ref}_{t})}[\mathbf{x}_{0}|\mathbf{x}_{t}^{ref}]$ is computed as an integral over $\mathbf{x}_{0}$, and it is expressed as the current state $\mathbf{x}_{t}^{ref}$ plus a term involving the score function $\nabla_{\mathbf{x}_{t}^{ref}}\log\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$, where $\sigma_{t}^{2}$ represents the noise variance. For notation simplicity, we denote $\mathbf{x}^{ref}_{t\rightarrow 0}$ as $\mathbb{E}_{\mathbf{x}_{0}\sim\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}^{ref}_{t})}[\mathbf{x}_{0}|\mathbf{x}_{t}^{ref}]$ without introducing ambiguity.
Therefore, essentially, we can obtain the expectation of $\mathbf{x}_{0}$ of $\mathbf{x}_{t}^{ref}$ (i.e., $\mathbf{x}^{ref}_{t\rightarrow 0}$) at arbitrary intermediate timesteps. Our idea is to replace the $\mathbf{x}_{0}^{ref}$ as adopted in the baseline with $\mathbf{x}_{t\rightarrow 0}^{ref}$, which eliminates the needs for fully simulation of generation process. Specifically, we can truncate the whole denoising process into
[…]
Further more, by incorporating real data for consideration, we can further truncate the left part by replacing the denoising trajectory from $\mathbf{x}_{T}^{ref}$ to $\mathbf{x}_{t^{\prime}}^{ref}$ with the noise injection from real data, i.e.,
[…]
where $\mathbf{x}_{t^{\prime}}^{ref}$ is obtained by adding noise from real data following the forward process $\mathbb{P}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$. When $t^{\prime}=T$ and $t=1$, we have $\mathbf{x}_{t^{\prime}}^{ref}=\mathbf{x}_{T}^{ref}$ and $\mathbf{x}_{t\rightarrow 0}^{ref}=\mathbf{x}_{1\rightarrow 0}^{ref}=\mathbf{x}_{0}^{ref}$. In this case, truncated diffusion fine-tuning essentially reduces to the baseline method. Thus, truncated diffusion fine-tuning can be viewed as a generalization of the baseline method, as it eliminates the need for fully solving the denoising process numerically, leading to significant improvements in training efficiency.
Figure 5: Visual Comparison on CogVideoX. Prompts are set as “A pretty lady at the beach and smiling, well drawn face” and “jake gyllenhaall drinking an ice cold beer, painting by egon schiele”. The results below or on the right are generated with self-npo, while the results above or on the left are without it.
Solving the distribution discrepancy. It is worthy noting that $\mathbf{x}^{ref}_{t\rightarrow 0}$ indeed has a different distribution with $\mathbf{x}_{0}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})$. This is because $\mathbb{P}(\mathbf{x}_{t\rightarrow 0}^{ref}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})$ can be a very smooth distribution, and $\mathbf{x}_{t\rightarrow 0}^{ref}$ is the weighted average value of many potential $\mathbf{x}_{0}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})$. Therefore, directly adopting the vanilla diffusion fine-tuning on $\mathbf{x}_{t\rightarrow 0}^{ref}$ can not achieve equivalent effect as the baseline method. We solve this distribution discrepancy by adopting a different noise adding strategy. Specifically, instead of directly adding noise to $\mathbf{x}_{t\rightarrow 0}^{ref}$ to obtain $\mathbf{x}_{s}$, i.e.,
$\mathbf{x}_{s}=\alpha_{s}\mathbf{x}_{t\rightarrow 0}^{ref}+\sigma_{s}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})\,,$ (15)
we choose to obtain $\mathbf{x}_{s}$ by adding noise to $\mathbf{x}_{t}^{ref}$, i.e.,
$\mathbf{x}_{s}=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref}+\sqrt{\sigma_{s}^{2}-\sigma_{t}^{2}\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})\,.$ (16)
Then we optimize the diffusion model $\boldsymbol{f}_{\boldsymbol{\theta}}$ (corresponding to $\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{0}\mid\boldsymbol{c})$) following the standard diffusion training with $\mathbf{x}_{t\rightarrow 0}^{ref}$ as $\mathbf{x}_{0}$-prediction target, i.e.,
$\min\limits_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref}),\,\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\|\mathbf{x}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2},$ (17)
In the following, we theoretically show the reasonability of this operation from three aspects: 1) Distribution of $\mathbf{x}_{s}$. We show the the distribution of $\mathbf{x}_{s}$ is equivalent to the noise perturbed distribution of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})$ (See Theorem II.1). 2) Gradient equivalent optimization target. We show our optimization objective Equation 17 has an gradient equivalent learning objective, which is the expectation value of all $\mathbf{x}_{t\rightarrow 0}^{ref}$ (See Theorem II.2). 3) Equivalent optimization objective of diffusion models. We show the gradient equivalent learning objective as proved in 2) is equivalent to the learning objective of normal diffusion model (See Theorem II.3).
4 Experiments
4.1 Validation Setup
To better evaluate the performance of our method, we tested three baseline models: a) Stable Diffusion v1-5, a text-to-image generation model with approximately 860 million parameters, widely used for creating high-quality images from textual prompts in tasks like digital art and concept visualization; b) Stable Diffusion XL (SDXL), an advanced version with around 2.6 billion parameters, enabling higher-resolution image generation (e.g., 1024x1024) and improved compositional coherence; c) CogVideoX-5B, a video generation model with about 5 billion parameters, designed for tasks like video synthesis and frame interpolation, leveraging its larger capacity for temporal modeling. Therefore, our testing covers diffusion models of varying sizes and encompasses mainstream text-to-image and text-to-video tasks, ensuring a comprehensive and well-designed baseline.
4.2 Comparison
Figure 6: Quantitative performance comparison with stable diffusion XL based models. All metrics are tested with official weights.
Method PickScore HPSv2.1 ImageReward Aesthetic
SDXL 22.06 28.04 0.6246 6.114
SDXL + NPO 22.25 28.98 0.6831 6.136
SDXL + Self-NPO 22.26 28.24 0.6697 6.226
Diff.-DPO 22.57 29.76 0.8624 6.099
Diff.-DPO + NPO 22.69 30.48 0.9210 6.112
Diff.-DPO + Self-NPO 22.67 29.83 0.8784 6.179
Juggernaut 22.66 30.25 0.9778 6.021
Juggernaut + Self-NPO 22.77 30.56 0.9921 6.031
Figure 7: User study analysis.
Figure 8: Winning ratio comparison. Self-NPO works seamlessly with SD1.5, SDXL, and those optimized for preferences, consistently improving both their generation quality and alignment with human preferences.
Quantitative comparison. For text-to-image generation, we evaluate our method quantitatively by adhering to prior research and utilizing the ‘test_unique’ split from Pick-a-pic as the benchmark for testing (kirstain2023pick, ). We utilize PickScore (kirstain2023pick, ), HPSv2 (wu2023human, ), ImageReward (xu2024imagereward, ), and Laion-Aesthetic (schuhmann2022laion, ) as our evaluation metrics. The outcomes of this quantitative assessment are detailed in Tables 1 and 7. These tables reveal that integrating Self-NPO with the base model and its preference-optimized variants consistently improves the aesthetic quality of the generated images. Beyond presenting average scores, as shown in Fig. 8, we also compute the percentage of samples generated from the same prompt that attain a higher preference score. The results produced with Self-NPO markedly surpass those generated without it. For text-to-video generation, we follow the experimental setup of VADER prabhudesai2024video but use the more advanced text-to-video diffusion model, CogVideoX-5B, as the baseline. We evaluate performance improvements by employing various prompt sets that focus on different topics, specifically selecting the prompt sets provided by VADER. We report the quantitative evaluation and winning ratios in Table 5. In most cases, Self-NPO consistently enhances both visual quality and aesthetic scores.
Qualitative comparison. We present the generated results in different scenarios in Fig. 1, Fig. 3, Fig. 5, and Fig. 9, demonstrating that self-NPO consistently improves both the generation visual quality and fidelity.
User study. We evaluate generation quality across four areas: Color and Lighting, High-Frequency Details, Low-Frequency Composition, and Text-Image Alignment. In Color and Lighting, users assess whether the images feature natural, visually pleasing color schemes and lighting. For High-Frequency Details, users look at the sharpness and texture details, especially edges and fine elements. In Low-Frequency Composition, they focus on the overall structure and balance of the image. Finally, for Text-Image Alignment, users judge how well the generated image matches the input text. The user study uses prompts from Pickapic ‘validation_unique‘, with different models generating images based on the same random seed. Users choose between "No Preference," (Draws) "Self-NPO is better," (Wins) or "Self-NPO is worse" (Losses) for each pair of images. We collect responses from 9 volunteers, each evaluating 100 pairs of images generated by Dreamshaper, totaling 900 votes. Results, shown in Fig. 7, reveal that NPO significantly improves high-frequency details, enhances color and lighting preferences, and helps with composition. It also shows an improvement in text-image alignment.
4.3 Applications.
Plug-and-play. Our method is not only applicable to the original stable diffusion-based models and their fine-tuned versions optimized through preference optimization, but also extends directly to high-quality stylized models fine-tuned on private datasets. As shown in Fig. 9, when leveraging the Dreamshaper model alongside the self-NPO-optimized model as an unconditional predictor, we observe substantial improvements in both image quality and aesthetic appeal.
Figure 9: Applications of Self-NPO. Prompts: “A giant eagle monster art”, “Mario in prison”, and “Girl, pretty, flower, hair, smile”. Results below/right use Self-NPO; above/left without it.
Controllable generation. Moreover, our approach seamlessly integrates with controllable generation techniques. By incorporating controllable plugins like T2I-Adapter mou2024t2i , we significantly boost the model’s generation quality under layout-controlled conditions.
Unconditional generation. It is well-established that unconditional generation typically falls short of conditional generation in terms of quality. This disparity is even more pronounced in text-to-image diffusion models, where, without conditional input, the model often struggles to produce anything other than incoherent, low-quality images. However, with the self-NPO-enhanced model, we can generate relatively high-quality images even in the absence of conditional guidance.
Table 1: Quantitative performance comparison with stable diffusion v1-5 based models. means the metrics are copied from NPO papers. Other metrics are tested with official weights.
Method PickScore HPSv2.1 ImageReward Aesthetic
DDPO 21.06 24.91 0.0817 5.591
D3PO 20.76 23.97 -0.1235 5.527
Diff.-SPO 21.41 26.85 0.1738 5.946
SD-1.5 20.75 26.84 0.1064 5.539
SD-1.5 + NPO 21.26 27.36 0.2028 5.667
SD-1.5 + Self-NPO 21.00 27.04 0.2816 5.609
Diff.-DPO 21.12 25.93 0.2651 5.648
Diff.-DPO + NPO (reg$=500$) 21.58 27.60 0.3101 5.762
Diff.-DPO + NPO (reg$=1000$) 21.43 27.36 0.3472 5.773
Diff.-DPO + Self-NPO 21.34 27.78 0.4085 5.710
SePPO 21.51 28.45 0.5981 5.892
SePPO + Self NPO 21.73 30.28 0.6744 6.014
Diff.-DPO + NPO (reg$=500$) 21.58 27.60 0.3101 5.762
Diff.-DPO + NPO (reg$=1000$) 21.43 27.36 0.3472 5.773
Diff.-DPO + Self-NPO 21.34 27.78 0.4085 5.710
DreamShaper 21.85 28.85 0.6819 6.143
DreamShaper + NPO ($\alpha=1.0$) 22.30 30.13 0.7258 6.234
DreamShaper + NPO ($\alpha=0.6$) 22.39 29.92 0.6034 6.492
DreamShaper + Self-NPO ($\alpha=1.0$) 22.20 30.40 0.8038 6.196
DreamShaper + Self-NPO ($\alpha=0.9$) 22.36 30.39 0.7738 6.345
DreamShaper + Self-NPO ($\alpha=0.8$) 22.34 29.96 0.6548 6.562
5 Conclusions
In this paper, we explore existing diffusion model-based preference optimization techniques, including vanilla preference optimization and negative preference optimization. We observe that these methods require large amounts of explicit preference annotations or fragile reward model training, which involve costly manual data labeling and reward model training. To address this, we introduce Self-NPO (Negative Preference Optimization), a method that learns directly from data generated by the diffusion model itself, thus eliminating the need for explicit preference optimization annotations. Additionally, we propose Truncated Diffusion Fine-tuning, which reduces the dependency on full data generation simulations, significantly improving training efficiency. Extensive experimental results validate the effectiveness of Self-NPO.
Limitations: Since Self-NPO does not rely on explicit preference annotations, it might have a lower performance bound compared to NPO. However, it is important to highlight that by eliminating the need for explicit annotations, Self-NPO broadens the scope of preference optimization, particularly in domains where acquiring such annotations is challenging or impractical.
Appendix
[…]
Appendix I Related Works
This section provides an overview of prior research in diffusion-based preference optimization.
I.1 Preference datasets and reward models
Studies such as Pick-a-pic [21], ImageReward [61], and HPSv2 [57] have gathered image pairs produced by diffusion models using identical prompts and annotated them with human preferences. Laion-Aesthetic [41] involves human ratings of real images on a scale from 1 to 10. These datasets are used to train models that predict preference scores, establishing a groundwork for future advancements in optimizing human preferences within diffusion models.
I.2 Differentiable reward
Approaches like DRaFT [6], AlignProp [35], and ReFL [61] utilize pre-trained ImageReward models to evaluate generated images and adjust the generative models based on gradients from these differentiable reward models. While these methods are direct and efficient, they often suffer from reward leakage due to the limitations of reward models, such as generating overly saturated images to inflate scores.
I.3 Reinforcement learning
Techniques such as DDPO [3] and DPOK [11] interpret the diffusion denoising process as a Markov decision process and employ reinforcement learning algorithms to align with human preferences. Some studies [68] enhance performance by scaling up the training process, typically using PPO or its variants.
I.4 Direct preference optimization
Diffusion-DPO [52] introduces a training objective that allows for direct optimization on preference-labeled image pairs without simulation. D3PO [63] and SPO [24] merge reinforcement learning with direct preference optimization, eliminating the need for specific score values during training.
I.5 Direct preference optimization with demonstration data
Some approaches [65, 7, 15, 66] construct preference pairs by using high-quality demonstration data as positive examples and self-generated data as negative examples. However, while these methods share similarities with Self-NPO in their use of self-generated data, they rely heavily on offline, high-quality demonstration datasets, resembling naive supervised fine-tuning aimed at fitting the demonstration data. This dependency caps their performance at the quality of the demonstration data and limits their applicability to more advanced models. For instance, related works primarily demonstrate results on Stable Diffusion v1-5 rather than more powerful models like Stable Diffusion-XL. In contrast, Self-NPO requires no demonstration data. It operates by training a distribution-preserving "bad" model to enhance generation quality through classifier-free guidance. This simplicity and generality make Self-NPO a more flexible and scalable approach.
I.6 Diffusion-NPO
Diffusion-NPO [54] presents negative preference optimization (NPO), which trains models to produce outputs contrary to human preferences, enhancing performance by steering the model away from undesired outcomes using classifier-free guidance (CFG). While Self-NPO aims for similar negative preference alignment, it differs by not requiring explicit preference annotations, thus avoiding the costly process of manual data pairing.
I.7 Other works related to negative preference optimization
In addition to preference optimization in diffusion models, several unpublished studies and efforts in language models have explored or referenced concepts similar to negative preference, which we discuss here. Community efforts, such as  [56, 10, 31], employ textual inversion [12] to compress negative preference representations into negative-conditional inputs, creating negative text embeddings. However, these representations are complex and primarily suppress features tied to specific keywords, resulting in limited effectiveness. In language model research, works like [67] focus on negative preference optimization to unlearn undesirable data, such as harmful concepts or private information, by optimizing a negative forget loss.
Appendix II Proofs
Distribution of $\mathbf{x}_{s}$.
Theorem II.1.
Given $\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$, we define $\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})=\mathcal{N}(\mathbf{x}_{s};\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref},\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})$, we show the equivalence between $\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ and $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{\text{ref}}\mid\boldsymbol{c})$.
Proof.
We can rewrite $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$ as
$\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})=\sum_{\mathbf{x}_{0}^{ref}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\,.$ (18)
We can decompose the sampling process of $\mathbf{x}_{s}^{ref}$ through re-parameterization as following:
$\begin{split}\mathbf{x}_{0}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c}),\\ \boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}),\\ \mathbf{x}_{s}^{ref}=\alpha_{s}\mathbf{x}_{0}^{ref}+\sigma_{s}\boldsymbol{\epsilon}\,.\end{split}$ (19)
Similarly, we can rewrite $\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ as
$\begin{split}&\quad\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})\\ &=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\\ &=\sum_{\mathbf{x}_{0}^{ref}}\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\\ &=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\sum_{\mathbf{x}_{0}^{ref}}\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\,.\end{split}$ (20)
Therefore, we can decompose the sampling process of
$\begin{split}\mathbf{x}_{0}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c}),\\ \boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}),\\ \mathbf{x}_{t}^{ref}=\alpha_{t}\mathbf{x}_{0}^{ref}+\sigma_{t}\boldsymbol{\epsilon}^{\prime},\\ \boldsymbol{\epsilon}^{*}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})\,,\\ \mathbf{x}_{s}=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref}+\sqrt{(\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})}\boldsymbol{\epsilon}^{*}\,.\end{split}$ (21)
We have
$\begin{split}\mathbf{x}_{s}&=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref}+\sqrt{(\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})}\boldsymbol{\epsilon}^{*}\,\\ &=\alpha_{s}\mathbf{x}_{0}+\sqrt{(\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})}\boldsymbol{\epsilon}^{*}+\sigma_{t}\frac{\alpha_{s}}{\alpha_{t}}\boldsymbol{\epsilon}^{\prime}\\ &=\alpha_{s}\mathbf{x}_{0}+\sqrt{\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2}+\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2}}\boldsymbol{\epsilon}\\ &=\alpha_{s}\mathbf{x}_{0}+\sigma_{s}\boldsymbol{\epsilon}\end{split}$ (22)
Therefore, we have shown the equivalence between the $\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})$ and $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$. ∎
Gradient equivalent optimization target.
Theorem II.2.
Following the give conditions, Equation 17 has a gradient equivalent optimization target, i.e.,
$\min\limits_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2},\,,$ (23)
where
$\bar{\mathbf{x}}_{t}^{ref}=\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}$ (24)
Proof.
We first decompose the squared differences of the functions, which will allow us to express them in a form suitable for integration and expectation. Using the bilinearity of the Euclidean norm, we first decompose the squared norms:
$\begin{split}&\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\\ =&\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}\|_{2}^{2}+\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}-2\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\end{split}$ (25)
and
$\begin{split}&\|\mathbf{x}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\\ =&\|\mathbf{x}_{t\rightarrow 0}^{ref}\|_{2}^{2}+\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}-2\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\,.\end{split}$ (26)
Note that $\mathbf{x}_{t\rightarrow 0}^{ref}$ and $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}$, we can ignore them as non-differentiable constants.
Then we compute the expectation of $\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}\|_{2}^{2}$ under the distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$, specifically,
$\begin{split}&\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\|f_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\\ =&\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{s}}\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\mathrm{d}\mathbf{x}_{t}^{ref}\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\mathrm{d}\mathbf{x}_{s}\mathrm{d}\mathbf{x}_{t}^{ref}\\ =&\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref}),\,\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\,.\end{split}$ (27)
In the above, we expressed the marginal distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})$ as the product of the conditional distribution $\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})$ and prior distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$. Then we apply Fubini’s Theorem to interchange the order of integration, obtaining the expected equivalent objective.
Next, we compute the expectation of the inner product $\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle$ over the distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$. We have
$\begin{split}&\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\\ =&\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})\langle\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{s}}\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{t}^{ref}\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\int_{\mathbf{x}_{s}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref})\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{s}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref}),\,\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\,.\end{split}$ (28)
Therefore, we have shown the Equation 23 and Equation 17 are gradient equivalent optimization objectives. Therefore, by adopting Equation 17, we are actually letting $\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})$ fit $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}$ following the formulation of Equation 23. ∎
Equivalent optimization objective of diffusion models. Following Theorem II.2, we have shown the ground truth of the learning target of $\mathbf{x}_{s}^{ref}$ of Truncated Diffusion Fine-tuning is $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}$. Following Tweedie’s formula, we know the ground truth of the learning target of $\mathbf{x}_{s}^{ref}$ of vanilla diffusion models is $\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}$. In the following, we show the equivalence of the two learning targets.
Theorem II.3.
Following the given conditions with $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}$, we have
$\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}$ (29)
Proof.
According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$ is
$\mathbf{x}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\,.$ (30)
Then we have
$\begin{split}&\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{ref}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\\ &\times\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\times\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref},\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\times\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref},\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\int_{\mathbf{x}_{t}^{ref}}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref},\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}\mathrm{d}\mathbf{x}_{0}^{ref}\\ &=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\int_{\mathbf{x}_{t}^{ref}}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref},\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}\mathrm{d}\mathbf{x}_{0}^{ref}\\ &=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\\ &=\mathbb{E}_{\mathbf{x}_{0}^{ref}\sim\mathbb{P}(\mathbf{x}_{0}^{ref}|\mathbf{x}^{ref}_{s})}[\mathbf{x}_{0}^{ref}|\mathbf{x}_{s}^{ref}]\,,\end{split}$ (31)
where $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}^{ref}_{s},\boldsymbol{c})=\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}$.
Appendix III Implementations
III.1 Pseudo training code
In Algorithm 1, we present the pseudocode for Truncated Diffusion Fine-Tuning, which generalizes the baseline method and eliminates the need for fully solving the denoising process numerically, leading to significant improvements in training efficiency.
Algorithm 1 Truncated Diffusion Fine-tuning
  Input: dataset $\mathcal{D}$, pretrained model parameter $\boldsymbol{\theta}$, learning rate $\eta$, number of inference steps $N$, truncated simulation steps $K$, $k$ step ODE solver $\Phi(\cdot,\cdot,\cdot,k;\boldsymbol{\theta}^{-})$, $d(\cdot,\cdot)$, $\lambda(\cdot,\cdot)$, and $\mu$
  $\boldsymbol{\theta}^{-}\leftarrow\boldsymbol{\theta}$
  repeat
     Sample $(\mathbf{x},\boldsymbol{c})\sim\mathcal{D}$ and $n\sim\mathcal{U}\llbracket 1,N-1\rrbracket$
     Sample $\mathbf{x}_{t_{n+1}}\sim\mathcal{N}(\alpha_{t_{n+1}}\mathbf{x};\sigma_{t_{n+1}}^{2}\mathbf{I})$
     $k\leftarrow\min\{K,n\}$
     $\mathbf{x}_{t_{n+1-k}}\leftarrow\Phi(\mathbf{x}_{t_{n+1}},t_{n+1},\boldsymbol{c},k;\boldsymbol{\theta}^{-})$
     $\mathbf{x}_{t_{n+1-k}\rightarrow 0}=\mathbf{x}_{t_{n+1-k}}+\sigma_{t_{n+1-k}}^{2}\nabla_{\mathbf{x}_{t_{n+1-k}}}\log\mathbb{P}_{\boldsymbol{\theta}^{-}}(\mathbf{x}_{t_{n+1-k}}\mid\boldsymbol{c})$
     Sample $m\sim\mathcal{U}\llbracket n+1-k,N-1\rrbracket$
     Sample $\mathbf{x}_{t_{m+1}}\sim\mathcal{N}\left(\frac{\alpha_{t_{m+1}}}{\alpha_{t_{n+1-k}}}\mathbf{x}_{t_{n+1-k}},\sqrt{\sigma_{t_{m+1}}^{2}-\frac{\alpha_{t_{m+1}}^{2}}{\alpha_{t_{n+1-k}}^{2}}\sigma_{t_{n+1-k}^{2}}}\mathbf{I}\right)$
     $\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\theta}^{-})\leftarrow\lambda(t_{n+1-k},t_{m+1})d(\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{t_{m+1}},t_{m+1},\boldsymbol{c}),\mathbf{x}_{t_{n+1-k}\rightarrow 0})$
     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\theta}^{-})$
     $\boldsymbol{\theta}^{-}\leftarrow\operatorname{stopgrad}(\mu\boldsymbol{\theta}^{-}+(1-\mu)\boldsymbol{\theta}$)
  until finish
III.2 Implementation details
For the training of text-to-image models, including Stable Diffusion v1-5 and Stable Diffusion XL, we utilize a random subset of 500,000 images sampled from the LAION dataset. In the case of training the text-to-video model, CogVideoX-5B, we select 20,000 video clips from the OpenVid dataset at random [30]. In all training scenarios, the number of inference steps $N$ is set to 25, and the truncated simulation steps $K$ is set to 5. To ensure a gradual degradation of generation quality, we do not fully freeze the weights used in the simulation; instead, we apply the exponential moving average (EMA) weights for training, defined as $\boldsymbol{\theta}^{-}=\mu\boldsymbol{\theta}^{-}+(1-\mu)\boldsymbol{\theta}$. For simplicity, we set $\lambda(\cdot,\cdot):=1$. All models are trained using four A800 GPUs. For Stable Diffusion v1-5, we use a learning rate of $1\times 10^{-6}$, a batch size of 80, and conduct training for 1000 iterations, which requires approximately 30 minutes (2 A800 GPU hours). For Stable Diffusion XL, the learning rate is set to $1\times 10^{-7}$, the batch size to 40, and training runs for 200 iterations, taking around 0.6 hours (2.4 A800 GPU hours). For CogVideoX, the learning rate is $1\times 10^{-6}$, the batch size is 4, and training lasts for about 600 iterations, which takes approximately 4 hours (16 A800 GPU hours).
Appendix IV Discussion
IV.1 Corrupted ODE solving
Notice that, in order to achieve negative preference optimization, we typically aim to have the unconditional/negative-conditional models learn to generate lower-quality results while still maintaining some level of distribution consistency. We discovered that we can intentionally create these lower-quality outputs by disrupting the ODE Solver’s solving process to a certain extent, using a negative classifier-free guidance (CFG) value (with $\omega=-5$ by default). The reason for this is that CFG has a considerable impact on the generation quality in text-conditioned diffusion models. By lowering the CFG value or even setting it to a negative value, we can simulate the generation of lower-quality data. As illustrated in Fig. 10, this method notably enhances the final generation quality and human alignment while incurring the same training cost.
Figure 10: Influence of corrupted ODE solving. Red lines denote the change curves of mean scores. Green lines denote the change curves of the winning ratios. Blue lines denote the mean score of original Stable Diffusion XL without applying Self-NPO.
IV.2 Sample filtering
To augment negative preference optimization, we incorporate sample filtering alongside corrupted ODE solving. We use PickScore as a reward model to evaluate generated samples within each batch, assigning higher loss weights to samples with lower scores (indicating poorer quality) and normalizing weights within batches. All other training and inference hyperparameters remain consistent with the standard Self-NPO setup. The results, presented in Table 2, compare Stable Diffusion v1-5 without Self-NPO, Self-NPO, and Self-NPO with filtering, using Aesthetic Score (AES), Human Preference Score (HPS), ImageReward, and PickScore. Win-rates are computed relative to the baseline.
Table 2: Performance of Self-NPO with and without sample filtering on Stable Diffusion v1-5. Win-rates are computed against Stable Diffusion v1-5 without Self-NPO.
Method AES Mean AES Win-Rate HPS Mean HPS Win-Rate ImageReward Mean ImageReward Win-Rate PickScore Mean PickScore Win-Rate
SD v1-5 w/o Self-NPO 5.5393 25.01 0.1064 20.76
Self-NPO 5.6189 57.6% 27.02 88.2% 0.2539 65.4% 20.97 68.8%
Self-NPO + Filtering 5.6392 60.2% 27.22 89.4% 0.2630 65.4% 21.02 71.2%
Sample filtering enhances Self-NPO’s performance across all metrics, with improvements in AES (5.6392 vs. 5.6189), HPS (27.22 vs. 27.02), ImageReward (0.2630 vs. 0.2539), and PickScore (21.02 vs. 20.97), as well as higher win-rates in AES (60.2% vs. 57.6%), HPS (89.4% vs. 88.2%), and PickScore (71.2% vs. 68.8%). However, this approach reintroduces a dependency on an external reward model, which may complicate the optimization process and introduce potential biases. Consequently, we exclude sample filtering from our default configuration to maintain simplicity and robustness in Self-NPO’s implementation.
IV.3 Training iterations
A nontrivial issue in training self-NPO is determining the training duration. Unlike standard diffusion training, negative preference optimization should not be over-optimized, as this would cause the unconditional predictions to deviate too much from the conditional predictions, leading to performance degradation. Inspired by early stopping technique [37], to determine the training duration, we select a set of prompts that are not used for benchmark testing and monitor the changes in the generation quality during training. When the quality stops improving or worsens, we stop the training. We found that usually, just a few training iterations are enough to see a significant improvement in performance.
IV.4 Training efficiency comparison
We reported the training costs of Self-NPO in our submission (Supplementary III.2 Implementation Details). With the default truncated simulation steps set to 5, we trained Stable Diffusion v1-5 for 1,000 iterations with a batch size of 80. The overall training time was approximately 0.5 hours on 4 A800 GPUs (2 A100 GPU hours). For the baseline (i.e., full simulation, $K=25$) described in our paper, training for 1,000 iterations required approximately 2.6 hours on 4 A800 GPUs (10.4 A800 GPU hours), which is around 5 times longer than our default setting.
According to the official GitHub page of Diffusion-DPO, the official weights of Diffusion-DPO were trained with a batch size of 2,048 for 2,000 iterations, taking 24 hours on 16 A100 GPUs (384 A100 GPU hours, approximately 192 times longer than our default setting). Since DPO-based Diffusion-NPO can be trained by simply reversing the preference pair, it incurs the same training cost. However, when Diffusion-NPO is trained with LoRA, our tests show that the training time is reduced to 153.6 A800 GPU hours, approximately 76.8 times longer than our default setting.
Table 3: Training costs for different methods.
Method Training GPU Hours GPU Type
Diffusion-NPO Full weight 384 A100
Diffusion-NPO LoRA 153.6 A800
Baseline Full weight 10.4 A800
Ours Full weight 2 A800
Additionally, we observed that the baseline’s performance is inferior to our method despite its higher training cost. We report the performance comparison below:
Table 4: Performance comparison between our method and the baseline. The baseline, shown in Fig. 4, is our method without truncated diffusion finetuning.
Method AES AES HPS HPS ImageReward ImageReward PickScore PickScore
Mean Score Win-Rate Mean Score Win-Rate Mean Score Win-Rate Mean Score Win-Rate
Baseline 5.5422 40.2% 26.64 40.4% 0.2342 46.6% 20.78 29.8%
Ours 5.6189 59.8% 27.02 59.6% 0.2539 53.4% 20.97 70.2%
IV.5 Advantages of Self-NPO
The key advantage of Self-NPO lies in its simplicity and ability to operate without explicit preference annotations, as emphasized in its title and introduction. Unlike prior methods such as DR, RL, DPO, and Diffusion-NPO, which rely on costly manual collection of preference data pairs and the complex, often unstable process of reward model training, Self-NPO eliminates these dependencies. This makes it highly practical, especially in domains where acquiring or labeling preference data is expensive or impractical.
In essence, Self-NPO acts as a "free lunch" for diffusion models, requiring only the original diffusion model itself without additional resources. It learns directly from the model, bypassing the need for external inputs.
For instance, while preview work [65, 15, 66] uses self-generated data to improve quality, it still depends on carefully curated, high-quality real image datasets as demonstration data. Similarly, Diffusion-NPO [54] seeks to train models aligned with negative preferences but requires either a reward model or preference data pairs. In contrast, Self-NPO’s simplicity and generality make it a versatile solution for all classifier-free guidance-supported diffusion models. It does not rely on model weights trained with preference annotations; instead, it can use pretrained weights directly (validated on SD v1.5, SDXL, and CogVideoX) or weights fine-tuned on datasets reflecting personal high-quality preferences (validated with Dreamshaper on SD v1.5 and Juggernaut on SDXL). This enables Self-NPO to perform negative preference optimization without explicit annotations, broadening its applicability in challenging domains.
In practice, many researchers and users share preference-optimized or high-quality fine-tuned models on public platforms, often without providing preference pairs or curated datasets. Self-NPO offers a plug-and-play approach to further enhance these models’ performance, making it a highly accessible and effective method.
Appendix V Additional Results
V.1 Text-to-video generation
We evaluate performance improvements by employing various prompt sets that focus on different topics, specifically selecting the prompt sets provided by VADER. We report the quantitative evaluation and winning ratios in Table 5. In most cases, Self-NPO consistently enhances both visual quality and aesthetic scores.
Table 5: Quantitative performance comparison on text-to-video generation. All metrics are tested with official weights. Avg means the average score. Win means the average winning ratio to other methods. The ‘Official’ means we use the prompts provided on the official page of CogVideoX. The other prompts follows the
Method AES HPS Image Reward PickScore
Avg Win Avg Win Avg Win Avg Win
Animal
CogVideoX 4.9492 35.37% 24.86 41.67% 0.0955 42.18% 0.2095 32.14%
CogVideoX + Self-NPO 4.9871 64.63% 24.98 58.33% 0.1746 57.82% 0.2106 67.86%
Offical
CogVideoX 5.3601 41.50% 24.52 25.51% 0.4852 53.06% 0.2025 39.29%
CogVideoX + Self-NPO 5.3968 58.50% 25.15 74.49% 0.6909 46.94% 0.2032 60.71%
Fashion
CogVideoX 5.2335 46.86% 22.96 34.94% 0.2462 46.04% 0.2062 50.37%
CogVideoX + Self-NPO 5.2368 53.14% 23.62 65.06% 0.3102 53.96% 0.2061 49.63%
Human
CogVideoX 5.0926 42.12% 23.99 34.86% 0.4593 51.51% 0.2142 45.59%
CogVideoX + Self-NPO 5.1417 57.88% 24.34 65.14% 0.5075 48.49% 0.2151 54.41%
Instrument
CogVideoX 5.1932 50.12% 27.06 39.31% 1.0015 48.20% 0.2218 46.49%
CogVideoX + Self-NPO 5.1866 49.88% 27.45 60.69% 1.0535 51.80% 0.2226 53.51%
Average
CogVideoX 5.1657 43.19% 24.68 35.26% 0.4575 48.20% 0.2108 42.78%
CogVideoX + Self-NPO 5.1898 56.81% 25.11 64.74% 0.5473 51.80% 0.2115 57.22%
Table 6: Comparison of different configurations on the metrics. All values are averaged over multiple tests.
Method AES HPS Image Reward PickScore
Avg Win Avg Win Avg Win Avg Win
CFG 5 Resolution (1344, 768)
w/o Self-NPO 5.9204 26.20% 26.56 45.80% 0.1478 45.20% 21.57 37.80%
w/ Self-NPO 6.1335 73.80% 26.63 54.20% 0.2066 54.80% 21.76 62.20%
CFG 5 Resolution (768, 1344)
w/o Self-NPO 5.9327 29.00% 26.60 40.60% 0.5454 54.20% 21.86 41.40%
w/ Self-NPO 6.1369 71.00% 26.68 59.40% 0.4862 45.80% 21.98 58.60%
CFG 5 Resolution (1024, 1024)
w/o Self-NPO 6.1051 33.20% 28.02 43.60% 0.6163 47.00% 22.10 32.40%
w/ Self-NPO 6.2344 66.80% 28.11 56.40% 0.6227 53.00% 22.26 67.60%
CFG 6 Resolution (1024, 1024)
w/o Self-NPO 6.0882 33.80% 28.46 46.00% 0.6813 45.20% 22.16 31.80%
w/ Self-NPO 6.2203 66.20% 28.64 54.00% 0.7583 54.80% 22.34 68.20%
CFG 7 Resolution (1024, 1024)
w/o Self-NPO 6.0949 33.60% 28.69 44.60% 0.7231 44.80% 22.16 32.20%
w/ Self-NPO 6.2037 66.40% 28.90 55.40% 0.7925 55.20% 22.37 67.80%
CFG 8 Resolution (1024, 1024)
w/o Self-NPO 6.0747 33.80% 28.89 44.80% 0.7655 47.20% 22.22 32.80%
w/ Self-NPO 6.2218 66.20% 28.96 55.20% 0.7755 52.80% 22.39 67.20%
CFG 9 Resolution (1024, 1024)
w/o Self-NPO 6.0691 34.20% 29.06 42.40% 0.8143 49.80% 22.22 31.80%
w/ Self-NPO 6.1952 65.80% 29.13 57.60% 0.7946 50.20% 22.41 68.20%
Unconditional
w/o Self-NPO 5.7107 31.25% 19.52 12.50% -0.1040 3.13% 19.50 28.13%
w/ Self-NPO 6.1102 68.75% 21.84 87.50% 0.0450 96.88% 19.77 71.88%
V.2 Effectiveness of Self-NPO on different inference scenarios
To assess the generalizability and practical effectiveness of the performance improvements facilitated by self-npo, we conducted extensive evaluations across diverse scenarios. The detailed results of these experiments are presented in Table 6.
  1. 1.
    Classifier-free guidance values. We examined the model’s performance with classifier-free guidance values set to 5, 6, 7, 8, and 9. In all tested configurations, the self-npo-enhanced model exhibited consistent and significant improvements across all four evaluation metrics, highlighting its robustness across different guidance intensities.
  2. 2.
    Generation resolutions. In addition to evaluating the model at the standard resolution of (1024, 1024), which is the resolution the self-npo model was specifically trained with, we also tested its performance on images with higher aspect ratios: longer (1344, 768) and wider (768, 1344). The results show that even when scaling up the generation resolutions, the self-npo model maintained or surpassed its original performance, demonstrating notable improvements across the majority of metrics in these non-standard configurations.
  3. 3.
    Unconditional generation. For the unconditional generation scenario, where no conditioning inputs are provided, the baseline Stable Diffusion XL model struggled to generate coherent results. In contrast, the self-npo-enhanced model delivered substantial improvements in generation quality, yielding marked performance gains across all evaluation metrics. This demonstrates the model’s capacity to enhance output even in challenging, unconstrained conditions.
V.3 Effectiveness of Self-NPO compared to training-free CFG-strengthening techniques
To further validate the efficacy of Self-NPO, we conducted experiments comparing it against training-free CFG-strengthening approaches, namely Autoguidance [19] and PAG [1].
Comparison with PAG on Stable Diffusion v1-5. We performed the following evaluations: 1) PAG on Stable Diffusion v1-5 vs. Naive CFG on Stable Diffusion v1-5. 2) PAG with CFG on Stable Diffusion v1-5 vs. Naive CFG on Stable Diffusion v1-5. 3) Stable Diffusion v1-5 as conditional predictors and Self-NPO-optimized Stable Diffusion v1-5 as unconditional predictors vs. Naive CFG on Stable Diffusion v1-5. Both PAG and PAG+CFG used the default hyperparameters from the official PAG GitHub repository. The results are presented in Table 7.
Table 7: Comparison of Self-NPO with PAG on Stable Diffusion v1-5.
Method AES Mean Win-Rate HPS Mean Win-Rate ImageReward Mean Win-Rate PickScore Mean Win-Rate
Naive CFG 5.5393 25.01 0.1064 20.76
PAG 5.3801 36.2% 21.82 17.0% -0.6493 21.4% 19.82 17.0%
PAG+CFG 5.5070 46.6% 24.35 39.2% -0.0799 39.4% 20.57 43.0%
Self-NPO 5.6189 57.6% 27.02 88.2% 0.2539 65.4% 20.97 68.8%
The results indicate that PAG and PAG+CFG, with their default hyperparameters, underperform compared to Naive CFG across most metrics. In contrast, Self-NPO demonstrates significant improvements across all evaluated metrics.
Comparison with Autoguidance on Stable Diffusion v1-5. We conducted the following experiments: 1) SePPO-optimized Stable Diffusion v1-5 as conditional predictors and original Stable Diffusion v1-5 as unconditional predictors vs. Naive CFG on SePPO-optimized Stable Diffusion v1-5 (baseline for Autoguidance). 2) SePPO-optimized Stable Diffusion v1-5 as conditional predictors and Self-NPO-optimized Stable Diffusion v1-5 as unconditional predictors vs. Naive CFG on SePPO-optimized Stable Diffusion v1-5. Note that the original Stable Diffusion v1-5 serves as an earlier version of the SePPO-optimized model, making the first setup the baseline for Autoguidance. The results are shown in Table 8.
Table 8: Comparison of Self-NPO with Autoguidance on Stable Diffusion v1-5.
Method AES Mean Win-Rate HPS Mean Win-Rate ImageReward Mean Win-Rate PickScore Mean Win-Rate
Baseline 5.8916 28.45 0.5918 21.51
Autoguidance 5.9912 61.0% 28.42 48.4% 0.4256 37.2% 20.87 20.0%
Self-NPO 6.0140 66.0% 30.28 87.2% 0.6744 60.4% 21.73 64.4%
Autoguidance shows a marginal improvement in Aesthetic Score but fails to improve other metrics. Self-NPO, however, consistently enhances performance across all metrics, demonstrating its superior effectiveness.
V.4 Performance of Self-NPO under different truncation values
To assess the influence of truncation values on the performance of Self-NPO, we conducted experiments using Stable Diffusion v1-5 with varying truncation steps ($K$) in the self-generation process. The truncation value controls the trade-off between simulation efficiency and the extent of self-generated content, with lower values prioritizing efficiency and higher values emphasizing comprehensive simulation. Our experiments evaluated $K\in\{1,2,5,10\}$, with $K=5$ selected as the default based on its balance of computational efficiency and performance across multiple metrics. The evaluation metrics include Aesthetic Score (AES), Human Preference Score (HPS), ImageReward, and PickScore, with win-rates computed relative to Stable Diffusion v1-5 without Self-NPO. The results are summarized in Table 9.
Table 9: Performance of Self-NPO on Stable Diffusion v1-5 with different truncation values ($K$). Win-rates are computed against Stable Diffusion v1-5 without Self-NPO.
Method AES Mean Win HPS Mean Win ImageReward Mean Win PickScore Mean Win
SD v1-5 w/o Self-NPO 5.5393 25.01 0.1064 20.76
$K=1$ 5.6613 65.6% 26.65 83.0% 0.2265 65.0% 21.01 69.4%
$K=2$ 5.6663 63.0% 26.83 85.4% 0.2481 63.8% 21.02 71.2%
$K=5$ 5.6189 57.6% 27.02 88.2% 0.2539 65.4% 20.97 68.8%
$K=10$ 5.6014 56.4% 27.44 91.0% 0.2445 63.8% 20.91 58.8%
The results demonstrate that Self-NPO consistently outperforms the baseline across all truncation values and metrics. Notably, $K=1$ and $K=2$ yield higher AES and PickScore win-rates, suggesting that minimal truncation can effectively enhance aesthetic quality and user preference. However, $K=5$ achieves the highest HPS mean score and a strong HPS win-rate (88.2%), indicating superior alignment with human preferences. Increasing $K$ to 10 further improves the HPS mean score (27.44) and win-rate (91.0%) but slightly reduces performance in AES and PickScore, likely due to increased computational complexity introducing minor artifacts. Given these findings, $K=5$ was selected as the default truncation value, as it provides an optimal balance between simulation efficiency and robust performance across diverse evaluation metrics.
V.5 Effectiveness of Self-NPO on domain specific traditional benchmark
To evaluate the generalizability of Self-NPO, we extended our experiments beyond large-scale latent diffusion models (Stable Diffusion v1-5, Stable Diffusion XL, and CogVideoX) to include domain-specific data. Specifically, we applied Self-NPO to the CIFAR-10 dataset, training a pixel diffusion model for class-conditional image generation following the implementation of EDM [18]. Unlike latent diffusion models, which operate on compressed representations, this setup directly processes raw pixel data, presenting a distinct challenge for generative modeling. We measured performance using two standard metrics: Fréchet Inception Distance (FID, lower is better) and Inception Score (higher is better), which assess the fidelity and diversity of generated images, respectively. The results are presented in Table 10.
Table 10: Performance of Self-NPO on CIFAR-10 compared to EDM, evaluated using FID and Inception Score.
Method FID Inception Score
EDM 1.8521 9.6375
EDM + Self-NPO 1.7472 10.0201
Improvement 0.1049 0.3826
The results show that Self-NPO significantly enhances the performance of the EDM baseline on CIFAR-10. The FID score improves by 0.1049, indicating that Self-NPO generates images with greater similarity to the ground-truth distribution. Similarly, the Inception Score increases by 0.3826, reflecting improved diversity and perceptual quality of the generated samples. These improvements demonstrate Self-NPO’s ability to enhance both the fidelity and variety of generated outputs in a domain-specific setting, where the model must capture fine-grained details directly from pixel data. The success on CIFAR-10, combined with our prior evaluations on large-scale models, underscores the versatility of Self-NPO across different data modalities and model architectures, making it a robust approach for improving diffusion-based generative modeling.
Appendix VI Societal Impact
VI.1 Positive Societal Impact
Our work on Self-NPO offers significant positive societal benefits by enhancing the accessibility and efficiency of diffusion models for visual content generation. By eliminating the need for costly manual preference annotations or reward model training, Self-NPO reduces the resource barriers for developers and creators, particularly in domains where preference data is scarce or expensive to acquire. This democratization fosters innovation and inclusivity, enabling a broader range of individuals and organizations to leverage high-quality image, video, and 3D content generation. Additionally, Self-NPO’s efficient tuning strategy minimizes computational costs, reducing energy consumption and contributing to more sustainable AI practices. The seamless integration with widely used models like SD1.5, SDXL, and CogVideoX further amplifies its impact, enhancing the quality and human preference alignment of generated content across diverse applications.
VI.2 Negative Societal Impact
Despite its benefits, Self-NPO, as a generative model optimization technique, carries potential risks. The ability to generate highly realistic visual content could be misused to produce misleading or harmful outputs, such as deepfakes or inappropriate imagery. Additionally, the automation of negative preference optimization without explicit human annotations may inadvertently amplify biases present in the pre-trained models, potentially leading to outputs that misalign with societal values or reinforce harmful stereotypes.
VI.3 Safeguards
To address these risks, we propose implementing robust safety measures. Self-NPO can leverage existing safety mechanisms in models like Stable Diffusion, which include internal content filters to detect and block harmful outputs. We also recommend integrating prompt-based harmful content detection to identify and halt generation processes triggered by problematic inputs. To mitigate potential biases, we advocate for regular audits of model outputs and the incorporation of fairness-aware fine-tuning techniques to ensure alignment with diverse societal values. By building on open-source frameworks, Self-NPO benefits from community-driven improvements and transparency, further enhancing its safety and reliability.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We have claimed in the abstract and introduction that we propose self-NPO, a negative preference optimization approach that learns exclusively from the model itself, thereby eliminating the need for manual data labeling or reward model training required by previous works.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of work in Sec. 5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: We provide a complete proof in Sec. II.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We have carefully introduced the framework architecture and training strategy in Sec. 3. Also, we enclose the pseudo training code, the detailed framework design in the Appendix. Eventually, We will release the code and model of this paper.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We promise we will release the code and model of this paper once accepted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify the training details in Sec. III.2 and test details in Sec. 4.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported, because it would be too computationally expensive. However, we indeed present plenty of qualitative results to demonstrate the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We have clarified the computational resources in Sec. III.2 and Sec. IV.4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The positive societal impacts are discussed in Sec. VI.1, and the negative societal impacts are discussed in Sec. VI.2.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Yes, our models are the image and video generators, which naturally have a risk of generating harmful content. Therefore, in Sec. VI.3, we introduce our solution to avoid harmful content. We will include the safeguards in our pipeline when releasing the code and model.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All the assets used in this paper are public datasets, allowing for research use.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not release new asserts in the submission. But we will release the code and model shortly. We will prepare documentation alongside the assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects. Guidelines
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not involve crowdsourcing, nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
-Self-NPO: Negative Preference Optimization of Diffusion Models by Simply Learning from Itself without Explicit Preference Annotations
+Self-NPO: Data-Free Diffusion Model Enhancement via Truncated Diffusion Fine-Tuning
 Abstract
@@
 While existing PO methods primarily concentrate on producing favorable outputs, they often overlook the significance of classifier-free guidance (CFG) in mitigating undesirable results.
-Diffusion-NPO addresses this gap by introducing negative preference optimization (NPO), training models to generate outputs opposite to human preferences and thereby steering them away from unfavorable outcomes.
-However, prior NPO approaches, including Diffusion-NPO, rely on costly and fragile procedures for obtaining explicit preference annotations (e.g., manual pairwise labeling or reward model training), limiting their practicality in domains where such data are scarce or difficult to acquire.
-In this work, we introduce Self-NPO, a Negative Preference Optimization approach that learns exclusively from the model itself, thereby eliminating the need for manual data labeling or reward model training.
-Moreover, our method is highly efficient and does not require exhaustive data sampling.
+Diffusion-NPO addresses this gap by introducing negative preference optimization (NPO), training models to generate outputs opposite to human preferences and thereby steering them away from unfavorable outcomes through CFG.
+However, prior NPO approaches rely on costly and fragile procedures for obtaining explicit preference annotations (e.g., manual pairwise labeling or reward model training), limiting their practicality in domains where such data are scarce or difficult to acquire.
+In this work, we propose Self-NPO, specifically truncated diffusion fine-tuning, a data-free approach of negative preference optimization by directly learning from the model itself, eliminating the need for manual data labeling or reward model training.
+This data-free approach is highly efficient (less than 1% training cost of Diffusion-NPO) and achieves comparable performance to Diffusion-NPO in a data-free manner.
 We demonstrate that Self-NPO integrates seamlessly into widely used diffusion models, including SD1.5, SDXL, and CogVideoX, as well as models already optimized for human preferences, consistently enhancing both their generation quality and alignment with human preferences.
-1 Introduction
-Over the past few years, diffusion models ho2020denoising ; song2020denoising ; song2020score have made significant strides in various visual generation tasks, including image karras2022elucidating ; rombach2022high ; wang2024rectified ; podell2023sdxl ; diffusionbeatgan ; meng2021sdedit ; teng2024dim ; ma2024storynizor ; li2024ecnet ; sun2024journeydb , video blattmann2023align ; mao2024osv ; shi2024motion ; makeavideo ; chen2024videocrafter2 ; bian2025gs ; NEURIPS2022_f7f47a73 ; wang2024animatelcm ; wang2025your, and 3D generation gao2024cat3d ; poole2022dreamfusion ; li2025connecting ; yan2025consistent ; sun2024genca ; lai2025unleashing.
+Code is available at https://github.com/G-U-N/Diffusion-NPO.
+Figure 1: Motivation of Self-NPO.
+Preference Optimization (PO) optimizes the preference aligned model by training on preference annotations.
+Negative Preference Optimization (NPO) optimizes the negative preference aligned model by training on switched preference annotations.
+Self-NPO optimizes the negative preference aligned model by truncated diffusion fine-tuning (TDFT) on truncated self-generated data.
+Introduction
+Figure 2: Visual Comparison.
+Prompts from left to right, from top to bottom are “galaxy cat”, “cat as a mafia boss”, and “A real picture of an Canguru with sunglasses at the sunset holding a stop sign”.
+Over the past few years, diffusion models (ho2020denoising; song2020denoising; song2020score) have made significant strides in various visual generation tasks, including image (karras2022elucidating; rombach2022high; wang2024rectified; podell2023sdxl; diffusionbeatgan; meng2021sdedit; teng2024dim; ma2024storynizor; li2024ecnet; sun2024journeydb), video (blattmann2023align; mao2024osv; shi2024motion; makeavideo; chen2024videocrafter2; bian2025gs; NEURIPS2022_f7f47a73; wang2024animatelcm; wang2025your; wang2024phased), and 3D generation (gao2024cat3d; poole2022dreamfusion; li2025connecting; yan2025consistent; sun2024genca; lai2025unleashing).
 However, diffusion models trained on massive unfiltered data (e.g., image-text pairs) often generate results that do not align well with human preferences.
 The growing interest in aligning these models for human preference-aligned generation has led to the development of preference optimization (PO) methods.
-In general, current PO methods can be categorized into four types:
-Figure 1: Visual Comparison.
-Prompts from left to right, from top to bottom are “galaxy cat”, “cat as a mafia boss”, and “A real picture of an Canguru with sunglasses at the sunset holding a stop sign”.
-Differentiable Reward (DR): In this approach, images generated via iterative denoising are evaluated by a pre-trained reward model, and the diffusion models are subsequently refined through gradient-based backpropagation to align with the reward model xu2024imagereward ; prabhudesai2024video ; zhang2024large ; wu2023hpsv1 ; wu2023human ; wu2024deepreward .
-Reinforcement Learning (RL): Here, the denoising process in diffusion models is treated as a Markov decision process (MDP) puterman2014markov ; sutton2018reinforcement .
-Techniques like PPO and its variants are employed to optimize preferences schulman2017proximal .
-Images are generated and evaluated dynamically based on reward feedback, with the goal of maximizing the chances of producing images that yield higher rewards.
-Direct Preference Optimization (DPO): DPO streamlines training by reframing the reinforcement learning objective into a simulation-free framework rafailov2024direct ; wallace2024diffusion , eliminating the need for separate reward models or online image generation and evaluation.
-Instead, it relies on fine-tuning with existing datasets of paired preference annotations.
-Some works yuan2024self ; deng2024enhancing ; gu2024diffusion ; zhang2024seppo construct preference pairs by using high-quality demonstration data as positive examples and self-generated data as negative examples.
-We have more discussion with these works as they also mention the idea of self-generated data.
-Negative Preference Optimization (Diffusion-NPO): Despite previous efforts to make models generate human-aligned images, Diffusion-NPO wang2025diffusionnpo firstly raised an important question: How can a model know to avoid generating poor images if it only knows how to generate good ones without understanding what is bad?
-Diffusion-NPO introduced the concept of negative preference optimization (NPO), which trains models to generate outputs that oppose human preferences, effectively improving performance by guiding the model away from undesired results though the classifier-free guidance (CFG) ho2022classifier ; karras2024guiding ; Shen_2024_CVPR ; ahn2024self .
-However, we argue that a major limitation of existing methods including DR, RL, DPO, and Diffusion-NPO, is their reliance on explicit preference annotations, such as the expensive manual collection of preference data pairs kirstain2023pick ; wu2023hpsv2 and the challenging and fragile process of reward model training xu2024imagereward ; wu2023hpsv2 ; schuhmann2022laion .
+In general, current PO methods can be categorized into four types: ($\mathrm{i}$) Differentiable Reward (DR) evaluates images from iterative denoising using a pre-trained reward model, refining diffusion models through gradient-based backpropagation to align with the reward model (xu2024imagereward; prabhudesai2024video; zhang2024large; wu2023hpsv1; wu2023human; wu2024deepreward; clark2023directly).
+($\mathrm{ii}$) Reinforcement Learning (RL) models the denoising process as a Markov decision process, using techniques like PPO to optimize preferences by dynamically generating and evaluating images to maximize rewards (puterman2014markov; sutton2018reinforcement; schulman2017proximal).
+($\mathrm{iii}$) Direct Preference Optimization (DPO) streamlines training by fine-tuning with paired preference datasets, eliminating separate reward models or online evaluation, often using high-quality data as positive examples and self-generated data as negatives (rafailov2024direct; wallace2024diffusion; yuan2024self; deng2024enhancing; gu2024diffusion; zhang2024seppo).
+($\mathrm{iv}$) Negative Preference Optimization (Diffusion-NPO) trains models to avoid poor outputs by opposing human preferences, improving performance through classifier-free guidance (CFG) to steer away from undesired results (wang2025diffusionnpo; ho2022classifier; karras2024guiding; Shen_2024_CVPR; ahn2024self).
+However, we argue that a major limitation of existing methods is their reliance on data sources, ie, explicit preference annotations, such as the expensive manual collection of preference data pairs (kirstain2023pick; wu2023hpsv2) and the challenging and fragile process of reward model training (xu2024imagereward; wu2023hpsv2; schuhmann2022laion).
 This heavy dependency on preference annotations makes these approaches less practical, particularly in domains where obtaining or labeling preference data is expensive or infeasible.
-In this work, we introduce Self-NPO, a novel approach for negative preference optimization (NPO) that can be done in an explicit preference annotation-free manner.
+In this work, we introduce Self-NPO, specifically, truncated diffusion fine-tuning (TDFT), a novel approach for negative preference optimization (NPO) that can be done in a data-free manner.
 Specifically, our training objective is to apply negative preference optimization to a pre-trained model, making its predictions more aligned with the opposite direction of human preferences.
-This facilitates unconditional/negative-conditional outputs in CFG, thus reducing the probability of generating results that do not align with human preferences.
+This facilitates unconditional/negative-conditional outputs in classifier-free guidance, thus reducing the probability of generating results that do not align with human preferences.
 To implement Self-NPO, we have three key observations:
@@
     To avoid altering the variance of the epsilon predictions, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{neg}}(\mathbf{x}_{t},t,\boldsymbol{c^{\prime}})$ should maintain a certain level of correlation with $\boldsymbol{\epsilon}_{\boldsymbol{\theta}_{pos}}(\mathbf{x}_{t},t,\boldsymbol{c})$.
-For example, for two completely independent Gaussian noises, the above operation would result in a variance of $12\omega^{2}+2\omega+$ 1.
+For example, for two completely independent Gaussian noises, the above operation would result in a variance of $2\omega^{2}+2\omega+1$.
   2. 2.
@@
 We do not need to fully execute iterative diffusion generation process to implement self-learning.
-We propose a novel tuning strategy, which allows us to fine tune diffusion models on the partially excuted iterative diffusion generation results.
+We propose a novel tuning strategy, termed as truncated diffusion fine-tuning, which allows us to fine tune diffusion models on the partially excuted iterative diffusion generation results.
 This avoids the huge generation costs of preparing data for self-learning.
-We demonstrate that Self-NPO can be seamlessly integrated with existing models, such as SD1.5 rombach2022high , SDXL podell2023sdxl , and CogVideoX yang2024cogvideox , as well as models previously fine-tuned through preference optimization, consistently improving their alignment with human preferences.
-Figure 2: Motivation of Self-NPO.
-Preference Optimization (PO) optimizes the preference aligned model by training on preference annotations.
-Negative Preference Optimization (NPO) optimizes the negative preference aligned model by training on switched preference annotations.
-Self-NPO optimizes the negative preference aligned model by truncated diffusion fine-tuning on truncated self-generated data.
-2 Preliminary
+We demonstrate that Self-NPO can be seamlessly integrated with existing models, such as SD1.5 (rombach2022high), SDXL (podell2023sdxl), and CogVideoX (yang2024cogvideox), as well as models previously fine-tuned through preference optimization, consistently improving their alignment with human preferences.
+Preliminary
 Preliminary of CFG.
 CFG has become a necessary and important technique for improving generation quality and text alignment of diffusion models.
-For convenience, we focus our discussion on the general formal of diffusion models, i.e., $\boldsymbol{x}_{t}=\alpha_{t}\boldsymbol{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ kingma2021variational .
-Suppose we learn a score estimator from a epsilon prediction neural network $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c},t)$, and we have $\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)=-\frac{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},t)}{\sigma_{t}}$.
+For convenience, we focus our discussion on the general formal of diffusion models, i.e., $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$ (kingma2021variational).
+Suppose we learn a score estimator from a epsilon prediction neural network $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\boldsymbol{c},t)$, and we have $\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t)=-\frac{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},t)}{\sigma_{t}}$.
 The sample prediction at timestep $t$ of the score estimator is formulated as
-$\hat{\boldsymbol{x}}_{0}=\frac{1}{\alpha_{t}}(\boldsymbol{x}_{t}+\sigma^{2}\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t))\,.$ (2)
-Applying the CFG is equivalent to add an additional score term, that is, we replace $\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)$ in Eq.
+$\hat{\mathbf{x}}_{0}=\frac{1}{\alpha_{t}}(\mathbf{x}_{t}+\sigma^{2}\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t))\,.$ (2)
+Applying the CFG is equivalent to add an additional score term, that is, we replace $\nabla_{\mathbf{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t)$ in Eq.
 2 with the following term,
-$\begin{split}\nabla_{\boldsymbol{x}_{t}}\log\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)+\nabla_{\boldsymbol{x}_{t}}\log\left[\frac{\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)}{\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c}^{\prime};t)}\right]^{\omega}\,,\end{split}$ (3)
+[…]
 where $\omega$ is to control the strength of CFG, $\boldsymbol{c}$ and $\boldsymbol{c}^{\prime}$ are conditional and unconditional/negative-conditional inputs, respectively.
-It is apparent that the generation will be pushed to high probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c};t)$ and relatively low probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t}|\boldsymbol{c}^{\prime};t)$.
+It is apparent that the generation will be pushed to high probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c};t)$ and relatively low probability region of $\mathbb{P}_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\boldsymbol{c}^{\prime};t)$.
 Write the above equation into the epsilon format, and then we have
-$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\theta}}=(\omega+1){\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c},t)}-\omega{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{x}_{t},\boldsymbol{c}^{\prime},t)}\,.$ (4)
-Diffusion-NPO wang2025diffusionnpo .
+$\boldsymbol{\epsilon}^{\omega}_{\boldsymbol{\theta}}=(\omega+1){\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\boldsymbol{c},t)}-\omega{\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\mathbf{x}_{t},\boldsymbol{c}^{\prime},t)}\,.$ (4)
+Diffusion-NPO (wang2025diffusionnpo).
 Diffusion models use classifier-free guidance (CFG) to improve generation quality by combining conditional and negative-conditional outputs to favor preferred results.
@@
 which ensures more stable and correlated outputs.
-3 Methodology
+Methodology
 Figure 3: Visual comparison on Stable Diffusion XL.
 Prompts: “A cute blue cat.”, “An attractive young woman rolling her eyes.”, and “Close-up of a head with smoke from ears, watching a smartphone, dynamic angle”.
-3.1 NPO via Self-Generated Data
-Reinforcement Learning with Human Feedback (RLHF) rafailov2024direct ; wallace2024diffusion ; griffith2013policy ; liu2025improving aims to optimize a conditional distribution $\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})$, where $\boldsymbol{c}\sim\mathcal{D}_{c}$, in order to maximize the expectation value of the associated reward model $R(\boldsymbol{c},\mathbf{x}_{0})$.
+Figure 4: Truncated diffusion fine-tuning (TDFT).
+Baseline method requires full simulation of generation process, introducing huge amount of generation costs.
+Naive truncation reduces the simulation costs but introduces distribution mismatch.
+TDFT reduces the simulation costs and maintains the target distribution.
+NPO via self-generated data
+Reinforcement Learning with Human Feedback (RLHF) (rafailov2024direct; wallace2024diffusion; griffith2013policy; liu2025improving) aims to optimize a conditional distribution $\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})$, where $\boldsymbol{c}\sim\mathcal{D}_{c}$, in order to maximize the expectation value of the associated reward model $R(\mathbf{x}_{0},\boldsymbol{c})$.
 Simultaneously, the optimization imposes a regularization term that penalizes deviations from a reference distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})$.
 Formally, the objective function is formulated as follows:
-$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R(\boldsymbol{c},\mathbf{x}_{0})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,,\end{split}$ (8)
+[…]
 where $\beta$ is a hyper-parameter that governs the balance between the reward maximization and the regularization effect imposed by the KL-divergence.
 The regularization is crucial because it prevents the model from straying too far from the distribution where the reward model is reliable, while also preserving generation diversity and preventing mode collapse to a single high-reward sample.
-Replacing $R(\mathbf{x}_{0},\boldsymbol{c})$ with $R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})$ yields the RLHF objective for Negative Preference Optimization (NPO):
-$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R_{\text{NPO}}(\boldsymbol{c},\mathbf{x}_{0})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,.\end{split}$ (9)
+Replacing $R(\mathbf{x}_{0},\boldsymbol{c})$ with $R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})$ yields the learning objective for Negative Preference Optimization (NPO):
+$\begin{split}\max_{\mathbb{P}_{\theta}}\mathbb{E}_{\boldsymbol{c}\sim\mathcal{D}_{c},\mathbf{x}_{0}\sim\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})}\left[R_{\text{NPO}}(\mathbf{x}_{0},\boldsymbol{c})\right]\\ -\beta D_{\text{KL}}\left[\mathbb{P}_{\theta}(\mathbf{x}_{0}\mid\boldsymbol{c})\parallel\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})\right]\,.\end{split}$ (9)
 where the first term encourages the samples generated from $\mathbb{P}_{\boldsymbol{\theta}}$ to have low reward scores, while the second term constrains the learned distribution to remain close to $\mathbb{P}_{\text{ref}}$.
@@
 Consequently, exploiting self-generated data provides a straightforward path toward distribution-regularized negative preference optimization.
-Figure 2 summarizes the core idea and situates it among related methods.
+Fig.
+1 summarizes the core idea and situates it among related methods.
 This approach is justified on two grounds: 1) Distribution preservation.
@@
 Imperfect optimization within the diffusion model itself can produce suboptimal results, such as blurry or noisy details, which leads to lower reward scores due to the misalignment between generated outputs and the desired target distribution.
-Indeed, we find it is helpful to intentionally corrupt the ODE sampling process to some extend, thus increasing the probability of reward reduction.
-We have more discussion in the Supplementary IV.1.
-Figure 4: Truncated Diffusion Fine-Tuning.
-Baseline method requires full simulation of generation process, introducing huge amount of generation costs.
-Naive Truncation reduces the simulation costs but introduces distribution mismatch.
-Truncated Diffusion Fine-tuning reduces the simulation costs and maintains the target distribution.
+Indeed, we find it is helpful to intentionally corrupt the ODE sampling process to some extent, thus increasing the probability of reward reduction.
+We have more discussion and empirical evidence in the supplementary material.
 However, despite its conceptual simplicity, fine-tuning a diffusion model on self-generated data can be computationally expensive, as it requires generating large amounts of training samples.
@@
 This method updates the model using partially generated diffusion samples, thereby avoiding the high costs of performing full diffusion simulations during training.
-3.2 Truncated Diffusion Fine-tuning
+Truncated diffusion fine-tuning
 Baseline: Fine-tuning on fully generated data.
@@
 Our method: Truncated Diffusion Fine-tuning.
-According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$, based on the conditional $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}_{t}^{ref})=\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}|\mathbf{x}_{0})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}$ according to Bayes’ rule, is related to the score of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ efron2011tweedie ; zhou2024score .
+According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$, based on the conditional $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}_{t}^{ref})=\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}|\mathbf{x}_{0})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}$ according to Bayes’ rule, is related to the score of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ (efron2011tweedie; zhou2024score).
 Specifically, the expectation is:
@@
 Specifically, we can truncate the whole denoising process into
-[…]
+$\mathbf{x}^{ref}_{T}\rightarrow\mathbf{x}^{ref}_{T-1}\rightarrow\dots\mathbf{x}^{ref}_{t}\xrightarrow{\text{Tweedie's formula}}\mathbf{x}_{t\rightarrow 0}^{ref}\,$, (13)
 Further more, by incorporating real data for consideration, we can further truncate the left part by replacing the denoising trajectory from $\mathbf{x}_{T}^{ref}$ to $\mathbf{x}_{t^{\prime}}^{ref}$ with the noise injection from real data, i.e.,
-[…]
+$\mathbf{x}^{ref}_{t^{\prime}}\rightarrow\mathbf{x}^{ref}_{t^{\prime}-1}\rightarrow\dots\mathbf{x}^{ref}_{t}\xrightarrow{\text{Tweedie's formula}}\mathbf{x}_{t\rightarrow 0}^{ref}\,$, (14)
 where $\mathbf{x}_{t^{\prime}}^{ref}$ is obtained by adding noise from real data following the forward process $\mathbb{P}(\mathbf{x}_{t}\mid\mathbf{x}_{0})$.
-When $t^{\prime}=T$ and $t=1$, we have $\mathbf{x}_{t^{\prime}}^{ref}=\mathbf{x}_{T}^{ref}$ and $\mathbf{x}_{t\rightarrow 0}^{ref}=\mathbf{x}_{1\rightarrow 0}^{ref}=\mathbf{x}_{0}^{ref}$.
+When $t^{\prime}=T$ and $t=1$, we have $\mathbf{x}_{t^{\prime}}^{ref}=\mathbf{x}_{T}^{ref}$ and $\mathbf{x}_{t\rightarrow 0}^{ref}=\mathbf{x}_{1\rightarrow 0}^{ref}=\mathbf{x}_{0}^{ref}\$.
 In this case, truncated diffusion fine-tuning essentially reduces to the baseline method.
 Thus, truncated diffusion fine-tuning can be viewed as a generalization of the baseline method, as it eliminates the need for fully solving the denoising process numerically, leading to significant improvements in training efficiency.
-Figure 5: Visual Comparison on CogVideoX.
-Prompts are set as “A pretty lady at the beach and smiling, well drawn face” and “jake gyllenhaall drinking an ice cold beer, painting by egon schiele”.
-The results below or on the right are generated with self-npo, while the results above or on the left are without it.
 Solving the distribution discrepancy.
@@
 $\min\limits_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref}),\,\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\|\mathbf{x}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2},$ (17)
-In the following, we theoretically show the reasonability of this operation from three aspects: 1) Distribution of $\mathbf{x}_{s}$.
-We show the the distribution of $\mathbf{x}_{s}$ is equivalent to the noise perturbed distribution of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})$ (See Theorem II.1).
-2) Gradient equivalent optimization target.
-We show our optimization objective Equation 17 has an gradient equivalent learning objective, which is the expectation value of all $\mathbf{x}_{t\rightarrow 0}^{ref}$ (See Theorem II.2).
-3) Equivalent optimization objective of diffusion models.
-We show the gradient equivalent learning objective as proved in 2) is equivalent to the learning objective of normal diffusion model (See Theorem II.3).
-4 Experiments
-4.1 Validation Setup
-To better evaluate the performance of our method, we tested three baseline models: a) Stable Diffusion v1-5, a text-to-image generation model with approximately 860 million parameters, widely used for creating high-quality images from textual prompts in tasks like digital art and concept visualization; b) Stable Diffusion XL (SDXL), an advanced version with around 2.6 billion parameters, enabling higher-resolution image generation (e.g., 1024x1024) and improved compositional coherence; c) CogVideoX-5B, a video generation model with about 5 billion parameters, designed for tasks like video synthesis and frame interpolation, leveraging its larger capacity for temporal modeling.
+Theoretical grounding
+We theoretically show the reasonability of proposed TDFT from three aspects:
+  1. 1.
+    Distribution of $\mathbf{x}_{s}$.
+We show that the distribution of $\mathbf{x}_{s}$ is equivalent to the noise-perturbed distribution of $\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{\text{ref}}\mid\boldsymbol{c})$ (See Thm.
+1).
+  2. 2.
+    Gradient equivalent optimization target.
+We show that our optimization objective, Eq.
+(17), has a gradient-equivalent learning objective, which is the expectation value of all $\mathbf{x}_{t\rightarrow 0}^{\text{ref}}$ (See Thm.
+2).
+  3. 3.
+    Equivalent optimization objective of diffusion models.
+We show that the gradient-equivalent learning objective, as proved in 2), is equivalent to the learning objective of a standard diffusion model (See Thm.
+3).
+Proof Sketch: Due to the limited space, we encourage readers to review the detailed proofs in Thm.
+1, Thm.
+2, and Thm.
+3.
+These proofs rigorously establish the equivalence of distributions, optimization objectives, and their connections to diffusion models, forming the foundation for understanding the validity and effectiveness of our proposed approach.
+Experiments
+Validation setup
+To better evaluate the performance of our method, we tested three baseline models: a) Stable Diffusion v1-5, a text-to-image generation model with approximately 860 million parameters, widely used for creating high-quality images from textual prompts in tasks like digital art and concept visualization; b) Stable Diffusion XL (SDXL), an advanced version with around 2.6 billion parameters, enabling higher-resolution image generation (e.g., 1024x1024) and improved compositional coherence; c) CogVideoX-5B, a video generation model with about 5 billion parameters, designed for tasks like video synthesis and frame interpolation, leveraging its capacity for temporal modeling.
 Therefore, our testing covers diffusion models of varying sizes and encompasses mainstream text-to-image and text-to-video tasks, ensuring a comprehensive and well-designed baseline.
-4.2 Comparison
+Comparison
 Figure 5: Quantitative performance comparison with stable diffusion XL based models. All metrics are tested with official weights.
-Method PickScore HPSv2.1 ImageReward Aesthetic
+Method P.S. HPS. I.R. AES.
 SDXL 22.06 28.04 0.6246 6.114
@@
 Quantitative comparison.
-For text-to-image generation, we evaluate our method quantitatively by adhering to prior research and utilizing the ‘test_unique’ split from Pick-a-pic as the benchmark for testing (kirstain2023pick, ).
-We utilize PickScore (kirstain2023pick, ), HPSv2 (wu2023human, ), ImageReward (xu2024imagereward, ), and Laion-Aesthetic (schuhmann2022laion, ) as our evaluation metrics.
-The outcomes of this quantitative assessment are detailed in Tables 1 and 7.
+For text-to-image generation, we evaluate our method quantitatively by adhering to prior research and utilizing the ‘test_unique’ split from Pick-a-pic as the benchmark for testing (kirstain2023pick).
+We utilize PickScore (kirstain2023pick) (P.S.), HPSv2.1 (wu2023human) (HPS.), ImageReward (xu2024imagereward) (I.R.), and Laion-Aesthetic (schuhmann2022laion) (AES.) as our evaluation metrics.
+The outcomes of this quantitative assessment are detailed in Tab.
+1 and Tab.
+5.
 These tables reveal that integrating Self-NPO with the base model and its preference-optimized variants consistently improves the aesthetic quality of the generated images.
 Beyond presenting average scores, as shown in Fig.
-8, we also compute the percentage of samples generated from the same prompt that attain a higher preference score.
+7, we also compute the percentage of samples generated from the same prompt that attain a higher preference score.
 The results produced with Self-NPO markedly surpass those generated without it.
-For text-to-video generation, we follow the experimental setup of VADER prabhudesai2024video but use the more advanced text-to-video diffusion model, CogVideoX-5B, as the baseline.
-We evaluate performance improvements by employing various prompt sets that focus on different topics, specifically selecting the prompt sets provided by VADER.
-We report the quantitative evaluation and winning ratios in Table 5.
-In most cases, Self-NPO consistently enhances both visual quality and aesthetic scores.
+For text-to-video generation, we provide detailed experimental comparisons and demonstrations in the supplementary material, where our method achieves clear performance improvements on the majority of metrics.
 Qualitative comparison.
 We present the generated results in different scenarios in Fig.
-1, Fig.
-3, Fig.
-5, and Fig.
-9, demonstrating that self-NPO consistently improves both the generation visual quality and fidelity.
+2, Fig.
+3, and Fig.
+8, demonstrating that self-NPO consistently improves both the generation visual quality and fidelity.
 User study.
@@
 Finally, for text-image alignment, users judge how well the generated image matches the input text.
-The user study uses prompts from Pickapic ‘validation_unique‘, with different models generating images based on the same random seed.
-Users choose between "No Preference," (Draws) "Self-NPO is better," (Wins) or "Self-NPO is worse" (Losses) for each pair of images.
+The user study uses prompts from Pickapic validation_unique dataset, with different models generating images based on the same random seed.
+Users choose between “No Preference” (Draws), “Self-NPO is better” (Wins), or “Self-NPO is worse” (Losses) for each pair of images.
 We collect responses from 9 volunteers, each evaluating 100 pairs of images generated by Dreamshaper, totaling 900 votes.
 Results, shown in Fig.
-7, reveal that NPO significantly improves high-frequency details, enhances color and lighting preferences, and helps with composition.
+6, reveal that NPO significantly improves high-frequency details, enhances color and lighting preferences, and helps with composition.
 It also shows an improvement in text-image alignment.
-4.3 Applications.
+Applications
 Plug-and-play.
@@
 As shown in Fig.
-9, when leveraging the Dreamshaper model alongside the self-NPO-optimized model as an unconditional predictor, we observe substantial improvements in both image quality and aesthetic appeal.
+8, when leveraging the Dreamshaper model alongside the self-NPO-optimized model as an unconditional predictor, we observe substantial improvements in both image quality and aesthetic appeal.
 Figure 8: Applications of Self-NPO. Prompts: “A giant eagle monster art”, “Mario in prison”, and “Girl, pretty, flower, hair, smile”. Results below/right use Self-NPO; above/left without it.
@@
 Moreover, our approach seamlessly integrates with controllable generation techniques.
-By incorporating controllable plugins like T2I-Adapter mou2024t2i , we significantly boost the model’s generation quality under layout-controlled conditions.
+By incorporating controllable plugins like T2I-Adapter (mou2024t2i), we significantly boost the model’s generation quality under layout-controlled conditions.
 Unconditional generation.
@@
 However, with the self-NPO-enhanced model, we can generate relatively high-quality images even in the absence of conditional guidance.
-Table 1: Quantitative performance comparison with stable diffusion v1-5 based models. means the metrics are copied from NPO papers. Other metrics are tested with official weights.
-Method PickScore HPSv2.1 ImageReward Aesthetic
+Table 1: Quantitative performance comparison with stable diffusion v1-5 based models. means the metrics are copied from Diffusion-NPO. Other metrics are tested with official weights.
+Method P.S. HPS. I.R. AES.
 DDPO 21.06 24.91 0.0817 5.591
@@
 Diff.-DPO 21.12 25.93 0.2651 5.648
-Diff.-DPO + NPO (reg$=500$) 21.58 27.60 0.3101 5.762
-Diff.-DPO + NPO (reg$=1000$) 21.43 27.36 0.3472 5.773
+Diff.-DPO + NPO (reg=500=500) 21.58 27.60 0.3101 5.762
+Diff.-DPO + NPO (reg=1000=1000) 21.43 27.36 0.3472 5.773
 Diff.-DPO + Self-NPO 21.34 27.78 0.4085 5.710
 SePPO 21.51 28.45 0.5981 5.892
-SePPO + Self NPO 21.73 30.28 0.6744 6.014
-Diff.-DPO + NPO (reg$=500$) 21.58 27.60 0.3101 5.762
-Diff.-DPO + NPO (reg$=1000$) 21.43 27.36 0.3472 5.773
-Diff.-DPO + Self-NPO 21.34 27.78 0.4085 5.710
+SePPO + Self-NPO 21.73 30.28 0.6744 6.014
 DreamShaper 21.85 28.85 0.6819 6.143
@@
 DreamShaper + Self-NPO ($\alpha=0.8$) 22.34 29.96 0.6548 6.562
-5 Conclusions
-In this paper, we explore existing diffusion model-based preference optimization techniques, including vanilla preference optimization and negative preference optimization.
-We observe that these methods require large amounts of explicit preference annotations or fragile reward model training, which involve costly manual data labeling and reward model training.
-To address this, we introduce Self-NPO (Negative Preference Optimization), a method that learns directly from data generated by the diffusion model itself, thus eliminating the need for explicit preference optimization annotations.
-Additionally, we propose Truncated Diffusion Fine-tuning, which reduces the dependency on full data generation simulations, significantly improving training efficiency.
-Extensive experimental results validate the effectiveness of Self-NPO.
-Limitations: Since Self-NPO does not rely on explicit preference annotations, it might have a lower performance bound compared to NPO.
-However, it is important to highlight that by eliminating the need for explicit annotations, Self-NPO broadens the scope of preference optimization, particularly in domains where acquiring such annotations is challenging or impractical.
-Appendix
-[…]
-[…]
-Appendix I Related Works
-This section provides an overview of prior research in diffusion-based preference optimization.
-I.1 Preference datasets and reward models
-Studies such as Pick-a-pic [21], ImageReward [61], and HPSv2 [57] have gathered image pairs produced by diffusion models using identical prompts and annotated them with human preferences.
-Laion-Aesthetic [41] involves human ratings of real images on a scale from 1 to 10.
-These datasets are used to train models that predict preference scores, establishing a groundwork for future advancements in optimizing human preferences within diffusion models.
-I.2 Differentiable reward
-Approaches like DRaFT [6], AlignProp [35], and ReFL [61] utilize pre-trained ImageReward models to evaluate generated images and adjust the generative models based on gradients from these differentiable reward models.
-While these methods are direct and efficient, they often suffer from reward leakage due to the limitations of reward models, such as generating overly saturated images to inflate scores.
-I.3 Reinforcement learning
-Techniques such as DDPO [3] and DPOK [11] interpret the diffusion denoising process as a Markov decision process and employ reinforcement learning algorithms to align with human preferences.
-Some studies [68] enhance performance by scaling up the training process, typically using PPO or its variants.
-I.4 Direct preference optimization
-Diffusion-DPO [52] introduces a training objective that allows for direct optimization on preference-labeled image pairs without simulation.
-D3PO [63] and SPO [24] merge reinforcement learning with direct preference optimization, eliminating the need for specific score values during training.
-I.5 Direct preference optimization with demonstration data
-Some approaches [65, 7, 15, 66] construct preference pairs by using high-quality demonstration data as positive examples and self-generated data as negative examples.
-However, while these methods share similarities with Self-NPO in their use of self-generated data, they rely heavily on offline, high-quality demonstration datasets, resembling naive supervised fine-tuning aimed at fitting the demonstration data.
-This dependency caps their performance at the quality of the demonstration data and limits their applicability to more advanced models.
-For instance, related works primarily demonstrate results on Stable Diffusion v1-5 rather than more powerful models like Stable Diffusion-XL.
-In contrast, Self-NPO requires no demonstration data.
-It operates by training a distribution-preserving "bad" model to enhance generation quality through classifier-free guidance.
-This simplicity and generality make Self-NPO a more flexible and scalable approach.
-I.6 Diffusion-NPO
-Diffusion-NPO [54] presents negative preference optimization (NPO), which trains models to produce outputs contrary to human preferences, enhancing performance by steering the model away from undesired outcomes using classifier-free guidance (CFG).
-While Self-NPO aims for similar negative preference alignment, it differs by not requiring explicit preference annotations, thus avoiding the costly process of manual data pairing.
-I.7 Other works related to negative preference optimization
-In addition to preference optimization in diffusion models, several unpublished studies and efforts in language models have explored or referenced concepts similar to negative preference, which we discuss here.
-Community efforts, such as  [56, 10, 31], employ textual inversion [12] to compress negative preference representations into negative-conditional inputs, creating negative text embeddings.
-However, these representations are complex and primarily suppress features tied to specific keywords, resulting in limited effectiveness.
-In language model research, works like [67] focus on negative preference optimization to unlearn undesirable data, such as harmful concepts or private information, by optimizing a negative forget loss.
-Appendix II Proofs
-Distribution of $\mathbf{x}_{s}$.
-Theorem II.1.
-Given $\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$, we define $\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})=\mathcal{N}(\mathbf{x}_{s};\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref},\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})$, we show the equivalence between $\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ and $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{\text{ref}}\mid\boldsymbol{c})$.
-Proof.
-We can rewrite $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$ as
-$\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})=\sum_{\mathbf{x}_{0}^{ref}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\,.$ (18)
-We can decompose the sampling process of $\mathbf{x}_{s}^{ref}$ through re-parameterization as following:
-$\begin{split}\mathbf{x}_{0}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c}),\\ \boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}),\\ \mathbf{x}_{s}^{ref}=\alpha_{s}\mathbf{x}_{0}^{ref}+\sigma_{s}\boldsymbol{\epsilon}\,.\end{split}$ (19)
-Similarly, we can rewrite $\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$ as
-$\begin{split}&\quad\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})\\ &=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\\ &=\sum_{\mathbf{x}_{0}^{ref}}\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\\ &=\sum_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\sum_{\mathbf{x}_{0}^{ref}}\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\,.\end{split}$ (20)
-Therefore, we can decompose the sampling process of
-$\begin{split}\mathbf{x}_{0}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c}),\\ \boldsymbol{\epsilon}^{\prime}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}),\\ \mathbf{x}_{t}^{ref}=\alpha_{t}\mathbf{x}_{0}^{ref}+\sigma_{t}\boldsymbol{\epsilon}^{\prime},\\ \boldsymbol{\epsilon}^{*}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})\,,\\ \mathbf{x}_{s}=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref}+\sqrt{(\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})}\boldsymbol{\epsilon}^{*}\,.\end{split}$ (21)
-We have
-$\begin{split}\mathbf{x}_{s}&=\frac{\alpha_{s}}{\alpha_{t}}\mathbf{x}_{t}^{ref}+\sqrt{(\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})}\boldsymbol{\epsilon}^{*}\,\\ &=\alpha_{s}\mathbf{x}_{0}+\sqrt{(\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2})}\boldsymbol{\epsilon}^{*}+\sigma_{t}\frac{\alpha_{s}}{\alpha_{t}}\boldsymbol{\epsilon}^{\prime}\\ &=\alpha_{s}\mathbf{x}_{0}+\sqrt{\sigma_{s}^{2}-\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2}+\frac{\alpha_{s}^{2}}{\alpha_{t}^{2}}\sigma_{t}^{2}}\boldsymbol{\epsilon}\\ &=\alpha_{s}\mathbf{x}_{0}+\sigma_{s}\boldsymbol{\epsilon}\end{split}$ (22)
-Therefore, we have shown the equivalence between the $\mathbb{Q}(\mathbf{x}_{s}\mid\boldsymbol{c})$ and $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$. ∎
-Gradient equivalent optimization target.
-Theorem II.2.
-Following the give conditions, Equation 17 has a gradient equivalent optimization target, i.e.,
-$\min\limits_{\boldsymbol{\theta}}\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2},\,,$ (23)
-where
-$\bar{\mathbf{x}}_{t}^{ref}=\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}$ (24)
-Proof.
-We first decompose the squared differences of the functions, which will allow us to express them in a form suitable for integration and expectation.
-Using the bilinearity of the Euclidean norm, we first decompose the squared norms:
-$\begin{split}&\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\\ =&\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}\|_{2}^{2}+\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}-2\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\end{split}$ (25)
-and
-$\begin{split}&\|\mathbf{x}_{t\rightarrow 0}^{ref}-\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\\ =&\|\mathbf{x}_{t\rightarrow 0}^{ref}\|_{2}^{2}+\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}-2\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\,.\end{split}$ (26)
-Note that $\mathbf{x}_{t\rightarrow 0}^{ref}$ and $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}$, we can ignore them as non-differentiable constants.
-Then we compute the expectation of $\|\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}\|_{2}^{2}$ under the distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$, specifically,
-$\begin{split}&\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\|f_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\\ =&\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{s}}\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\mathrm{d}\mathbf{x}_{t}^{ref}\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\mathrm{d}\mathbf{x}_{s}\mathrm{d}\mathbf{x}_{t}^{ref}\\ =&\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref}),\,\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\|\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\|_{2}^{2}\,.\end{split}$ (27)
-In the above, we expressed the marginal distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})$ as the product of the conditional distribution $\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref},\boldsymbol{c})$ and prior distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})$.
-Then we apply Fubini’s Theorem to interchange the order of integration, obtaining the expected equivalent objective.
-Next, we compute the expectation of the inner product $\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle$ over the distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})$.
-We have
-$\begin{split}&\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\\ =&\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})\langle\bar{\mathbf{x}}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{s}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})\langle\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{s}}\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{t}^{ref}\mathrm{d}\mathbf{x}_{s}\\ =&\int_{\mathbf{x}_{t}^{ref}}\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})\int_{\mathbf{x}_{s}}\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref})\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\mathrm{d}\mathbf{x}_{s}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\mathbb{E}_{\mathbf{x}_{s}\sim\mathbb{P}(\mathbf{x}_{s}\mid\mathbf{x}_{t}^{ref}),\,\mathbf{x}_{t}^{ref}\sim\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\langle\mathbf{x}_{t\rightarrow 0}^{ref},\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})\rangle\,.\end{split}$ (28)
-Therefore, we have shown the Equation 23 and Equation 17 are gradient equivalent optimization objectives.
-Therefore, by adopting Equation 17, we are actually letting $\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{s},s,\boldsymbol{c})$ fit $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}$ following the formulation of Equation 23.
-∎
-Equivalent optimization objective of diffusion models.
-Following Theorem II.2, we have shown the ground truth of the learning target of $\mathbf{x}_{s}^{ref}$ of Truncated Diffusion Fine-tuning is $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}$.
-Following Tweedie’s formula, we know the ground truth of the learning target of $\mathbf{x}_{s}^{ref}$ of vanilla diffusion models is $\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}$.
-In the following, we show the equivalence of the two learning targets.
-Theorem II.3.
-Following the given conditions with $\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{t}^{ref}}\mathbf{x}_{t\rightarrow 0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}$, we have
-$\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}$ (29)
-Proof.
-According to the Tweedie’s formula, for the semi-implicit distribution $\mathbb{P}_{\text{ref}}(\mathbf{x}^{ref}_{t}\mid\boldsymbol{c})$ induced by the reference diffusion model, the expected value of $\mathbf{x}_{0}$ given $\mathbf{x}_{t}^{ref}$ is
-$\mathbf{x}_{t\rightarrow 0}^{ref}=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\,.$ (30)
-Then we have
-$\begin{split}&\bar{\mathbf{x}}_{t\rightarrow 0}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{ref}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}\\ &\times\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}_{\text{ref}}(\mathbf{x}_{t}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref})\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\times\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{t}^{ref},\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}(\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\times\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{t}^{ref}}\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref},\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\mathrm{d}\mathbf{x}_{t}^{ref}\\ &=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\int_{\mathbf{x}_{t}^{ref}}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref},\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}\mathrm{d}\mathbf{x}_{0}^{ref}\\ &=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\int_{\mathbf{x}_{t}^{ref}}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref},\mathbf{x}_{t}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{t}^{ref}\mathrm{d}\mathbf{x}_{0}^{ref}\\ &=\int_{\mathbf{x}_{0}^{ref}}\mathbf{x}_{0}^{ref}\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}\mathrm{d}\mathbf{x}_{0}^{ref}\\ &=\mathbb{E}_{\mathbf{x}_{0}^{ref}\sim\mathbb{P}(\mathbf{x}_{0}^{ref}|\mathbf{x}^{ref}_{s})}[\mathbf{x}_{0}^{ref}|\mathbf{x}_{s}^{ref}]\,,\end{split}$ (31)
-where $\mathbb{P}(\mathbf{x}_{0}|\mathbf{x}^{ref}_{s},\boldsymbol{c})=\frac{\mathbb{P}(\mathbf{x}_{s}^{ref}\mid\mathbf{x}_{0}^{ref},\boldsymbol{c})\mathbb{P}_{\text{ref}}(\mathbf{x}_{0}^{ref}\mid\boldsymbol{c})}{\mathbb{P}_{\text{ref}}(\mathbf{x}_{s}^{ref}\mid\boldsymbol{c})}$.
-Appendix III Implementations
-III.1 Pseudo training code
-In Algorithm 1, we present the pseudocode for Truncated Diffusion Fine-Tuning, which generalizes the baseline method and eliminates the need for fully solving the denoising process numerically, leading to significant improvements in training efficiency.
-Algorithm 1 Truncated Diffusion Fine-tuning
-  Input: dataset $\mathcal{D}$, pretrained model parameter $\boldsymbol{\theta}$, learning rate $\eta$, number of inference steps $N$, truncated simulation steps $K$, $k$ step ODE solver $\Phi(\cdot,\cdot,\cdot,k;\boldsymbol{\theta}^{-})$, $d(\cdot,\cdot)$, $\lambda(\cdot,\cdot)$, and $\mu$
-  $\boldsymbol{\theta}^{-}\leftarrow\boldsymbol{\theta}$
-  repeat
-     Sample $(\mathbf{x},\boldsymbol{c})\sim\mathcal{D}$ and $n\sim\mathcal{U}\llbracket 1,N-1\rrbracket$
-     Sample $\mathbf{x}_{t_{n+1}}\sim\mathcal{N}(\alpha_{t_{n+1}}\mathbf{x};\sigma_{t_{n+1}}^{2}\mathbf{I})$
-     $k\leftarrow\min\{K,n\}$
-     $\mathbf{x}_{t_{n+1-k}}\leftarrow\Phi(\mathbf{x}_{t_{n+1}},t_{n+1},\boldsymbol{c},k;\boldsymbol{\theta}^{-})$
-     $\mathbf{x}_{t_{n+1-k}\rightarrow 0}=\mathbf{x}_{t_{n+1-k}}+\sigma_{t_{n+1-k}}^{2}\nabla_{\mathbf{x}_{t_{n+1-k}}}\log\mathbb{P}_{\boldsymbol{\theta}^{-}}(\mathbf{x}_{t_{n+1-k}}\mid\boldsymbol{c})$
-     Sample $m\sim\mathcal{U}\llbracket n+1-k,N-1\rrbracket$
-     Sample $\mathbf{x}_{t_{m+1}}\sim\mathcal{N}\left(\frac{\alpha_{t_{m+1}}}{\alpha_{t_{n+1-k}}}\mathbf{x}_{t_{n+1-k}},\sqrt{\sigma_{t_{m+1}}^{2}-\frac{\alpha_{t_{m+1}}^{2}}{\alpha_{t_{n+1-k}}^{2}}\sigma_{t_{n+1-k}^{2}}}\mathbf{I}\right)$
-     $\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\theta}^{-})\leftarrow\lambda(t_{n+1-k},t_{m+1})d(\boldsymbol{f}_{\boldsymbol{\theta}}(\mathbf{x}_{t_{m+1}},t_{m+1},\boldsymbol{c}),\mathbf{x}_{t_{n+1-k}\rightarrow 0})$
-     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}(\boldsymbol{\theta},\boldsymbol{\theta}^{-})$
-     $\boldsymbol{\theta}^{-}\leftarrow\operatorname{stopgrad}(\mu\boldsymbol{\theta}^{-}+(1-\mu)\boldsymbol{\theta}$)
-  until finish
-III.2 Implementation details
-For the training of text-to-image models, including Stable Diffusion v1-5 and Stable Diffusion XL, we utilize a random subset of 500,000 images sampled from the LAION dataset.
-In the case of training the text-to-video model, CogVideoX-5B, we select 20,000 video clips from the OpenVid dataset at random [30].
-In all training scenarios, the number of inference steps $N$ is set to 25, and the truncated simulation steps $K$ is set to 5.
-To ensure a gradual degradation of generation quality, we do not fully freeze the weights used in the simulation; instead, we apply the exponential moving average (EMA) weights for training, defined as $\boldsymbol{\theta}^{-}=\mu\boldsymbol{\theta}^{-}+(1-\mu)\boldsymbol{\theta}$.
-For simplicity, we set $\lambda(\cdot,\cdot):=1$.
-All models are trained using four A800 GPUs.
-For Stable Diffusion v1-5, we use a learning rate of $1\times 10^{-6}$, a batch size of 80, and conduct training for 1000 iterations, which requires approximately 30 minutes (2 A800 GPU hours).
-For Stable Diffusion XL, the learning rate is set to $1\times 10^{-7}$, the batch size to 40, and training runs for 200 iterations, taking around 0.6 hours (2.4 A800 GPU hours).
-For CogVideoX, the learning rate is $1\times 10^{-6}$, the batch size is 4, and training lasts for about 600 iterations, which takes approximately 4 hours (16 A800 GPU hours).
-Appendix IV Discussion
-IV.1 Corrupted ODE solving
-Notice that, in order to achieve negative preference optimization, we typically aim to have the unconditional/negative-conditional models learn to generate lower-quality results while still maintaining some level of distribution consistency.
-We discovered that we can intentionally create these lower-quality outputs by disrupting the ODE Solver’s solving process to a certain extent, using a negative classifier-free guidance (CFG) value (with $\omega=-5$ by default).
-The reason for this is that CFG has a considerable impact on the generation quality in text-conditioned diffusion models.
-By lowering the CFG value or even setting it to a negative value, we can simulate the generation of lower-quality data.
-As illustrated in Fig.
-10, this method notably enhances the final generation quality and human alignment while incurring the same training cost.
-Figure 10: Influence of corrupted ODE solving.
-Red lines denote the change curves of mean scores.
-Green lines denote the change curves of the winning ratios.
-Blue lines denote the mean score of original Stable Diffusion XL without applying Self-NPO.
-IV.2 Sample filtering
-To augment negative preference optimization, we incorporate sample filtering alongside corrupted ODE solving.
-We use PickScore as a reward model to evaluate generated samples within each batch, assigning higher loss weights to samples with lower scores (indicating poorer quality) and normalizing weights within batches.
-All other training and inference hyperparameters remain consistent with the standard Self-NPO setup.
-The results, presented in Table 2, compare Stable Diffusion v1-5 without Self-NPO, Self-NPO, and Self-NPO with filtering, using Aesthetic Score (AES), Human Preference Score (HPS), ImageReward, and PickScore.
-Win-rates are computed relative to the baseline.
-Table 2: Performance of Self-NPO with and without sample filtering on Stable Diffusion v1-5. Win-rates are computed against Stable Diffusion v1-5 without Self-NPO.
-Method AES Mean AES Win-Rate HPS Mean HPS Win-Rate ImageReward Mean ImageReward Win-Rate PickScore Mean PickScore Win-Rate
-SD v1-5 w/o Self-NPO 5.5393 25.01 0.1064 20.76
-Self-NPO 5.6189 57.6% 27.02 88.2% 0.2539 65.4% 20.97 68.8%
-Self-NPO + Filtering 5.6392 60.2% 27.22 89.4% 0.2630 65.4% 21.02 71.2%
-Sample filtering enhances Self-NPO’s performance across all metrics, with improvements in AES (5.6392 vs.
-5.6189), HPS (27.22 vs.
-27.02), ImageReward (0.2630 vs.
-0.2539), and PickScore (21.02 vs.
-20.97), as well as higher win-rates in AES (60.2% vs.
-57.6%), HPS (89.4% vs.
-88.2%), and PickScore (71.2% vs.
-68.8%).
-However, this approach reintroduces a dependency on an external reward model, which may complicate the optimization process and introduce potential biases.
-Consequently, we exclude sample filtering from our default configuration to maintain simplicity and robustness in Self-NPO’s implementation.
-IV.3 Training iterations
-A nontrivial issue in training self-NPO is determining the training duration.
-Unlike standard diffusion training, negative preference optimization should not be over-optimized, as this would cause the unconditional predictions to deviate too much from the conditional predictions, leading to performance degradation.
-Inspired by early stopping technique [37], to determine the training duration, we select a set of prompts that are not used for benchmark testing and monitor the changes in the generation quality during training.
-When the quality stops improving or worsens, we stop the training.
-We found that usually, just a few training iterations are enough to see a significant improvement in performance.
-IV.4 Training efficiency comparison
-We reported the training costs of Self-NPO in our submission (Supplementary III.2 Implementation Details).
+Training efficiency comparison
+We report the training costs of Self-NPO.
 With the default truncated simulation steps set to 5, we trained Stable Diffusion v1-5 for 1,000 iterations with a batch size of 80.
@@
 Ours Full weight 2 A800
-Additionally, we observed that the baseline’s performance is inferior to our method despite its higher training cost. We report the performance comparison below:
-Table 4: Performance comparison between our method and the baseline. The baseline, shown in Fig. 4, is our method without truncated diffusion finetuning.
-Method AES AES HPS HPS ImageReward ImageReward PickScore PickScore
-Mean Score Win-Rate Mean Score Win-Rate Mean Score Win-Rate Mean Score Win-Rate
-Baseline 5.5422 40.2% 26.64 40.4% 0.2342 46.6% 20.78 29.8%
-Ours 5.6189 59.8% 27.02 59.6% 0.2539 53.4% 20.97 70.2%
-IV.5 Advantages of Self-NPO
-The key advantage of Self-NPO lies in its simplicity and ability to operate without explicit preference annotations, as emphasized in its title and introduction.
-Unlike prior methods such as DR, RL, DPO, and Diffusion-NPO, which rely on costly manual collection of preference data pairs and the complex, often unstable process of reward model training, Self-NPO eliminates these dependencies.
-This makes it highly practical, especially in domains where acquiring or labeling preference data is expensive or impractical.
-In essence, Self-NPO acts as a "free lunch" for diffusion models, requiring only the original diffusion model itself without additional resources.
-It learns directly from the model, bypassing the need for external inputs.
-For instance, while preview work [65, 15, 66] uses self-generated data to improve quality, it still depends on carefully curated, high-quality real image datasets as demonstration data.
-Similarly, Diffusion-NPO [54] seeks to train models aligned with negative preferences but requires either a reward model or preference data pairs.
-In contrast, Self-NPO’s simplicity and generality make it a versatile solution for all classifier-free guidance-supported diffusion models.
-It does not rely on model weights trained with preference annotations; instead, it can use pretrained weights directly (validated on SD v1.5, SDXL, and CogVideoX) or weights fine-tuned on datasets reflecting personal high-quality preferences (validated with Dreamshaper on SD v1.5 and Juggernaut on SDXL).
-This enables Self-NPO to perform negative preference optimization without explicit annotations, broadening its applicability in challenging domains.
-In practice, many researchers and users share preference-optimized or high-quality fine-tuned models on public platforms, often without providing preference pairs or curated datasets.
-Self-NPO offers a plug-and-play approach to further enhance these models’ performance, making it a highly accessible and effective method.
-Appendix V Additional Results
-V.1 Text-to-video generation
-We evaluate performance improvements by employing various prompt sets that focus on different topics, specifically selecting the prompt sets provided by VADER.
-We report the quantitative evaluation and winning ratios in Table 5.
-In most cases, Self-NPO consistently enhances both visual quality and aesthetic scores.
-Table 5: Quantitative performance comparison on text-to-video generation.
-All metrics are tested with official weights.
-Avg means the average score.
-Win means the average winning ratio to other methods.
-The ‘Official’ means we use the prompts provided on the official page of CogVideoX.
-The other prompts follows the
-Method AES HPS Image Reward PickScore
-Avg Win Avg Win Avg Win Avg Win
-Animal
-CogVideoX 4.9492 35.37% 24.86 41.67% 0.0955 42.18% 0.2095 32.14%
-CogVideoX + Self-NPO 4.9871 64.63% 24.98 58.33% 0.1746 57.82% 0.2106 67.86%
-Offical
-CogVideoX 5.3601 41.50% 24.52 25.51% 0.4852 53.06% 0.2025 39.29%
-CogVideoX + Self-NPO 5.3968 58.50% 25.15 74.49% 0.6909 46.94% 0.2032 60.71%
-Fashion
-CogVideoX 5.2335 46.86% 22.96 34.94% 0.2462 46.04% 0.2062 50.37%
-CogVideoX + Self-NPO 5.2368 53.14% 23.62 65.06% 0.3102 53.96% 0.2061 49.63%
-Human
-CogVideoX 5.0926 42.12% 23.99 34.86% 0.4593 51.51% 0.2142 45.59%
-CogVideoX + Self-NPO 5.1417 57.88% 24.34 65.14% 0.5075 48.49% 0.2151 54.41%
-Instrument
-CogVideoX 5.1932 50.12% 27.06 39.31% 1.0015 48.20% 0.2218 46.49%
-CogVideoX + Self-NPO 5.1866 49.88% 27.45 60.69% 1.0535 51.80% 0.2226 53.51%
-Average
-CogVideoX 5.1657 43.19% 24.68 35.26% 0.4575 48.20% 0.2108 42.78%
-CogVideoX + Self-NPO 5.1898 56.81% 25.11 64.74% 0.5473 51.80% 0.2115 57.22%
-Table 6: Comparison of different configurations on the metrics. All values are averaged over multiple tests.
-Method AES HPS Image Reward PickScore
-Avg Win Avg Win Avg Win Avg Win
-CFG 5 Resolution (1344, 768)
-w/o Self-NPO 5.9204 26.20% 26.56 45.80% 0.1478 45.20% 21.57 37.80%
-w/ Self-NPO 6.1335 73.80% 26.63 54.20% 0.2066 54.80% 21.76 62.20%
-CFG 5 Resolution (768, 1344)
-w/o Self-NPO 5.9327 29.00% 26.60 40.60% 0.5454 54.20% 21.86 41.40%
-w/ Self-NPO 6.1369 71.00% 26.68 59.40% 0.4862 45.80% 21.98 58.60%
-CFG 5 Resolution (1024, 1024)
-w/o Self-NPO 6.1051 33.20% 28.02 43.60% 0.6163 47.00% 22.10 32.40%
-w/ Self-NPO 6.2344 66.80% 28.11 56.40% 0.6227 53.00% 22.26 67.60%
-CFG 6 Resolution (1024, 1024)
-w/o Self-NPO 6.0882 33.80% 28.46 46.00% 0.6813 45.20% 22.16 31.80%
-w/ Self-NPO 6.2203 66.20% 28.64 54.00% 0.7583 54.80% 22.34 68.20%
-CFG 7 Resolution (1024, 1024)
-w/o Self-NPO 6.0949 33.60% 28.69 44.60% 0.7231 44.80% 22.16 32.20%
-w/ Self-NPO 6.2037 66.40% 28.90 55.40% 0.7925 55.20% 22.37 67.80%
-CFG 8 Resolution (1024, 1024)
-w/o Self-NPO 6.0747 33.80% 28.89 44.80% 0.7655 47.20% 22.22 32.80%
-w/ Self-NPO 6.2218 66.20% 28.96 55.20% 0.7755 52.80% 22.39 67.20%
-CFG 9 Resolution (1024, 1024)
-w/o Self-NPO 6.0691 34.20% 29.06 42.40% 0.8143 49.80% 22.22 31.80%
-w/ Self-NPO 6.1952 65.80% 29.13 57.60% 0.7946 50.20% 22.41 68.20%
-Unconditional
-w/o Self-NPO 5.7107 31.25% 19.52 12.50% -0.1040 3.13% 19.50 28.13%
-w/ Self-NPO 6.1102 68.75% 21.84 87.50% 0.0450 96.88% 19.77 71.88%
-V.2 Effectiveness of Self-NPO on different inference scenarios
-To assess the generalizability and practical effectiveness of the performance improvements facilitated by self-npo, we conducted extensive evaluations across diverse scenarios.
-The detailed results of these experiments are presented in Table 6.
-  1. 1.
-    Classifier-free guidance values.
-We examined the model’s performance with classifier-free guidance values set to 5, 6, 7, 8, and 9.
-In all tested configurations, the self-npo-enhanced model exhibited consistent and significant improvements across all four evaluation metrics, highlighting its robustness across different guidance intensities.
-  2. 2.
-    Generation resolutions.
-In addition to evaluating the model at the standard resolution of (1024, 1024), which is the resolution the self-npo model was specifically trained with, we also tested its performance on images with higher aspect ratios: longer (1344, 768) and wider (768, 1344).
-The results show that even when scaling up the generation resolutions, the self-npo model maintained or surpassed its original performance, demonstrating notable improvements across the majority of metrics in these non-standard configurations.
-  3. 3.
-    Unconditional generation.
-For the unconditional generation scenario, where no conditioning inputs are provided, the baseline Stable Diffusion XL model struggled to generate coherent results.
-In contrast, the self-npo-enhanced model delivered substantial improvements in generation quality, yielding marked performance gains across all evaluation metrics.
-This demonstrates the model’s capacity to enhance output even in challenging, unconstrained conditions.
-V.3 Effectiveness of Self-NPO compared to training-free CFG-strengthening techniques
-To further validate the efficacy of Self-NPO, we conducted experiments comparing it against training-free CFG-strengthening approaches, namely Autoguidance [19] and PAG [1].
-Comparison with PAG on Stable Diffusion v1-5.
-We performed the following evaluations: 1) PAG on Stable Diffusion v1-5 vs.
-Naive CFG on Stable Diffusion v1-5.
-2) PAG with CFG on Stable Diffusion v1-5 vs.
-Naive CFG on Stable Diffusion v1-5.
-3) Stable Diffusion v1-5 as conditional predictors and Self-NPO-optimized Stable Diffusion v1-5 as unconditional predictors vs.
-Naive CFG on Stable Diffusion v1-5.
-Both PAG and PAG+CFG used the default hyperparameters from the official PAG GitHub repository.
-The results are presented in Table 7.
-Table 7: Comparison of Self-NPO with PAG on Stable Diffusion v1-5.
-Method AES Mean Win-Rate HPS Mean Win-Rate ImageReward Mean Win-Rate PickScore Mean Win-Rate
-Naive CFG 5.5393 25.01 0.1064 20.76
-PAG 5.3801 36.2% 21.82 17.0% -0.6493 21.4% 19.82 17.0%
-PAG+CFG 5.5070 46.6% 24.35 39.2% -0.0799 39.4% 20.57 43.0%
-Self-NPO 5.6189 57.6% 27.02 88.2% 0.2539 65.4% 20.97 68.8%
-The results indicate that PAG and PAG+CFG, with their default hyperparameters, underperform compared to Naive CFG across most metrics.
-In contrast, Self-NPO demonstrates significant improvements across all evaluated metrics.
-Comparison with Autoguidance on Stable Diffusion v1-5.
-We conducted the following experiments: 1) SePPO-optimized Stable Diffusion v1-5 as conditional predictors and original Stable Diffusion v1-5 as unconditional predictors vs.
-Naive CFG on SePPO-optimized Stable Diffusion v1-5 (baseline for Autoguidance).
-2) SePPO-optimized Stable Diffusion v1-5 as conditional predictors and Self-NPO-optimized Stable Diffusion v1-5 as unconditional predictors vs.
-Naive CFG on SePPO-optimized Stable Diffusion v1-5.
-Note that the original Stable Diffusion v1-5 serves as an earlier version of the SePPO-optimized model, making the first setup the baseline for Autoguidance.
-The results are shown in Table 8.
-Table 8: Comparison of Self-NPO with Autoguidance on Stable Diffusion v1-5.
-Method AES Mean Win-Rate HPS Mean Win-Rate ImageReward Mean Win-Rate PickScore Mean Win-Rate
-Baseline 5.8916 28.45 0.5918 21.51
-Autoguidance 5.9912 61.0% 28.42 48.4% 0.4256 37.2% 20.87 20.0%
-Self-NPO 6.0140 66.0% 30.28 87.2% 0.6744 60.4% 21.73 64.4%
-Autoguidance shows a marginal improvement in Aesthetic Score but fails to improve other metrics.
-Self-NPO, however, consistently enhances performance across all metrics, demonstrating its superior effectiveness.
-V.4 Performance of Self-NPO under different truncation values
-To assess the influence of truncation values on the performance of Self-NPO, we conducted experiments using Stable Diffusion v1-5 with varying truncation steps ($K$) in the self-generation process.
-The truncation value controls the trade-off between simulation efficiency and the extent of self-generated content, with lower values prioritizing efficiency and higher values emphasizing comprehensive simulation.
-Our experiments evaluated $K\in\{1,2,5,10\}$, with $K=5$ selected as the default based on its balance of computational efficiency and performance across multiple metrics.
-The evaluation metrics include Aesthetic Score (AES), Human Preference Score (HPS), ImageReward, and PickScore, with win-rates computed relative to Stable Diffusion v1-5 without Self-NPO.
-The results are summarized in Table 9.
-Table 9: Performance of Self-NPO on Stable Diffusion v1-5 with different truncation values ($K$). Win-rates are computed against Stable Diffusion v1-5 without Self-NPO.
-Method AES Mean Win HPS Mean Win ImageReward Mean Win PickScore Mean Win
-SD v1-5 w/o Self-NPO 5.5393 25.01 0.1064 20.76
-$K=1$ 5.6613 65.6% 26.65 83.0% 0.2265 65.0% 21.01 69.4%
-$K=2$ 5.6663 63.0% 26.83 85.4% 0.2481 63.8% 21.02 71.2%
-$K=5$ 5.6189 57.6% 27.02 88.2% 0.2539 65.4% 20.97 68.8%
-$K=10$ 5.6014 56.4% 27.44 91.0% 0.2445 63.8% 20.91 58.8%
-The results demonstrate that Self-NPO consistently outperforms the baseline across all truncation values and metrics.
-Notably, $K=1$ and $K=2$ yield higher AES and PickScore win-rates, suggesting that minimal truncation can effectively enhance aesthetic quality and user preference.
-However, $K=5$ achieves the highest HPS mean score and a strong HPS win-rate (88.2%), indicating superior alignment with human preferences.
-Increasing $K$ to 10 further improves the HPS mean score (27.44) and win-rate (91.0%) but slightly reduces performance in AES and PickScore, likely due to increased computational complexity introducing minor artifacts.
-Given these findings, $K=5$ was selected as the default truncation value, as it provides an optimal balance between simulation efficiency and robust performance across diverse evaluation metrics.
-V.5 Effectiveness of Self-NPO on domain specific traditional benchmark
-To evaluate the generalizability of Self-NPO, we extended our experiments beyond large-scale latent diffusion models (Stable Diffusion v1-5, Stable Diffusion XL, and CogVideoX) to include domain-specific data.
-Specifically, we applied Self-NPO to the CIFAR-10 dataset, training a pixel diffusion model for class-conditional image generation following the implementation of EDM [18].
-Unlike latent diffusion models, which operate on compressed representations, this setup directly processes raw pixel data, presenting a distinct challenge for generative modeling.
-We measured performance using two standard metrics: Fréchet Inception Distance (FID, lower is better) and Inception Score (higher is better), which assess the fidelity and diversity of generated images, respectively.
-The results are presented in Table 10.
-Table 10: Performance of Self-NPO on CIFAR-10 compared to EDM, evaluated using FID and Inception Score.
-Method FID Inception Score
-EDM 1.8521 9.6375
-EDM + Self-NPO 1.7472 10.0201
-Improvement 0.1049 0.3826
-The results show that Self-NPO significantly enhances the performance of the EDM baseline on CIFAR-10.
-The FID score improves by 0.1049, indicating that Self-NPO generates images with greater similarity to the ground-truth distribution.
-Similarly, the Inception Score increases by 0.3826, reflecting improved diversity and perceptual quality of the generated samples.
-These improvements demonstrate Self-NPO’s ability to enhance both the fidelity and variety of generated outputs in a domain-specific setting, where the model must capture fine-grained details directly from pixel data.
-The success on CIFAR-10, combined with our prior evaluations on large-scale models, underscores the versatility of Self-NPO across different data modalities and model architectures, making it a robust approach for improving diffusion-based generative modeling.
-Appendix VI Societal Impact
-VI.1 Positive Societal Impact
-Our work on Self-NPO offers significant positive societal benefits by enhancing the accessibility and efficiency of diffusion models for visual content generation.
-By eliminating the need for costly manual preference annotations or reward model training, Self-NPO reduces the resource barriers for developers and creators, particularly in domains where preference data is scarce or expensive to acquire.
-This democratization fosters innovation and inclusivity, enabling a broader range of individuals and organizations to leverage high-quality image, video, and 3D content generation.
-Additionally, Self-NPO’s efficient tuning strategy minimizes computational costs, reducing energy consumption and contributing to more sustainable AI practices.
-The seamless integration with widely used models like SD1.5, SDXL, and CogVideoX further amplifies its impact, enhancing the quality and human preference alignment of generated content across diverse applications.
-VI.2 Negative Societal Impact
-Despite its benefits, Self-NPO, as a generative model optimization technique, carries potential risks.
-The ability to generate highly realistic visual content could be misused to produce misleading or harmful outputs, such as deepfakes or inappropriate imagery.
-Additionally, the automation of negative preference optimization without explicit human annotations may inadvertently amplify biases present in the pre-trained models, potentially leading to outputs that misalign with societal values or reinforce harmful stereotypes.
-VI.3 Safeguards
-To address these risks, we propose implementing robust safety measures.
-Self-NPO can leverage existing safety mechanisms in models like Stable Diffusion, which include internal content filters to detect and block harmful outputs.
-We also recommend integrating prompt-based harmful content detection to identify and halt generation processes triggered by problematic inputs.
-To mitigate potential biases, we advocate for regular audits of model outputs and the incorporation of fairness-aware fine-tuning techniques to ensure alignment with diverse societal values.
-By building on open-source frameworks, Self-NPO benefits from community-driven improvements and transparency, further enhancing its safety and reliability.
-NeurIPS Paper Checklist
-The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact.
-Do not remove the checklist: The papers not including the checklist will be desk rejected.
-The checklist should follow the references and follow the (optional) supplemental material.
-The checklist does NOT count towards the page limit.
-Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
-  • You should answer [Yes] , [No] , or [N/A] .
-  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
-  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
-The checklist answers are an integral part of your paper submission.
-They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers.
-You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
-The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation.
-While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used").
-In general, answering "[No] " or "[N/A] " is not grounds for rejection.
-While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate.
-All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix.
-If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
-IMPORTANT, please:
-  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
-  • Keep the checklist subsection headings, questions/answers and guidelines below.
-  • Do not modify the questions and only use the provided macros for your answers.
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4.
-Justification: We have claimed in the abstract and introduction that we propose self-NPO, a negative preference optimization approach that learns exclusively from the model itself, thereby eliminating the need for manual data labeling or reward model training required by previous works.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: We discuss the limitations of work in Sec. 5.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: We provide a complete proof in Sec. II.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19.
-Justification: We have carefully introduced the framework architecture and training strategy in Sec.
-3.
-Also, we enclose the pseudo training code, the detailed framework design in the Appendix.
-Eventually, We will release the code and model of this paper.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [No]
-  24. Justification: We promise we will release the code and model of this paper once accepted.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We specify the training details in Sec. III.2 and test details in Sec. 4.1.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [No]
-  34.
-Justification: Error bars are not reported, because it would be too computationally expensive.
-However, we indeed present plenty of qualitative results to demonstrate the statistical significance of the experiments.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We have clarified the computational resources in Sec. III.2 and Sec. IV.4.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research conducted in the paper conforms, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [Yes]
-  49. Justification: The positive societal impacts are discussed in Sec. VI.1, and the negative societal impacts are discussed in Sec. VI.2.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [Yes]
-  54.
-Justification: Yes, our models are the image and video generators, which naturally have a risk of generating harmful content.
-Therefore, in Sec.
-VI.3, we introduce our solution to avoid harmful content.
-We will include the safeguards in our pipeline when releasing the code and model.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: All the assets used in this paper are public datasets, allowing for research use.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: We do not release new asserts in the submission. But we will release the code and model shortly. We will prepare documentation alongside the assets.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: This paper does not involve crowdsourcing nor research with human subjects. Guidelines
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We do not involve crowdsourcing, nor research with human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Conclusions
+In this paper, we explore existing diffusion model-based preference optimization techniques, including vanilla preference optimization and negative preference optimization.
+We observe that these methods require large amounts of explicit preference annotations or fragile reward model training, which involve costly manual data labeling and reward model training.
+To address this, we introduce Self-NPO (Negative Preference Optimization), a method that learns directly from data generated by the diffusion model itself, thus eliminating the need for explicit preference optimization annotations.
+Additionally, we propose Truncated Diffusion Fine-tuning, which reduces the dependency on full data generation simulations, significantly improving training efficiency.
+Extensive experimental results validate the effectiveness of Self-NPO.
+Limitations: Since Self-NPO does not rely on explicit preference annotations, it might have a lower performance bound compared to NPO.
+However, it is important to highlight that by eliminating the need for explicit annotations, Self-NPO broadens the scope of preference optimization, particularly in domains where acquiring such annotations is challenging or impractical.
+Acknowledgement.
+This study was supported in part by National Key R&D Program of China Project 2022ZD0161100, in part by the Centre for Perceptual and Interactive Intelligence, a CUHK-led InnoCentre under the InnoHK initiative of the Innovation and Technology Commission of the Hong Kong Special Administrative Region Government, in part by NSFC-RGC Project N_CUHK498/24, and in part by Guangdong Basic and Applied Basic Research Foundation (No.
+2023B1515130008, XW).